### v1

# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

Abstract
Evolutionary multitasking (EMT) has been attracting much attention over the past years. It aims to handle multiple optimization tasks simultaneously within limited computing resources assisted by inter-task knowledge transfer techniques. Numerous multitask evolutionary algorithms (MTEAs) for solving multitask optimization (MTO) problems have been proposed in the EMT field, but there lacks a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To address this issue, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. It incorporates more than 30 MTEAs, more than 150 MTO problem cases with real-world applications, and more than 10 performance metrics. Moreover, for comparing MTEAs with traditional evolutionary algorithms, we modified more than 30 popular single-task evolutionary algorithms to be able to solve MTO problems in MToP. MToP is a user-friendly tool with a graphical user interface that makes it easy to analyze results, export data, and plot schematics. More importantly, MToP is extensible, allowing users to develop new algorithms and define new problems. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
Index Terms:
Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm.
I Introduction
Evolutionary computation (EC) is inspired by natural evolution and has grown rapidly due to effectiveness and efficiency. With its robustness and ease of use, EC algorithms, also known as evolutionary algorithms (EAs), have achieved great success in solving black-box optimization problems. Designing specific EAs for different complex black-box optimization problems, such as constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6], has also been studied extensively. In recent years, with the increase in computational demands and the development of cloud computing, the use of EAs to handle a large number of optimization tasks simultaneously has gained much focus. This requirement is called evolutionary multitasking (EMT) [7]. A multitask optimization (MTO) problem of EMT with $K$ optimization tasks aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks. It can be formulated as follows:
$\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,F_{K}(\bm{x}_{K})\Big{]},\\ \textit{s.t. }\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$ (1)
where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task. For multi-objective optimization task, its $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$ task. EMT has been successfully applied to engineering scheduling [5], feature selection in machine learning [8], anomaly detection [9], point cloud registration [10], and reinforcement learning [11].
In order to solve multiple optimization tasks faster and better concurrently, researchers have attempted to exploit the similarity among tasks to empower EAs with inter-task knowledge extraction and transfer [12]. Through knowledge transfer, EAs are then able to fully utilize implicit parallelism to obtain better solutions for multiple tasks with fewer computational resources [13]. The first attempt for EMT is the multifactorial EA [14], which proposes an implicit knowledge representation through random mating among optimization tasks. After that, a large number of EAs for solving MTO problems have been proposed. These developed multitask EAs (MTEAs) employ either a multifactorial framework [15, 16] that utilizes a single population for multiple tasks or a multi-population framework [17, 18] that assign separate populations for multiple tasks. In addition, for decision space mapping among different tasks, various techniques in MTEAs such as unified search space [14, 16], autoencoding [19, 20], affine transformation [21, 22], and adversarial generative model [23] have been successively proposed. Due to the significant influence of the amount of knowledge transfer for solving MTO problems, MTEAs with knowledge transfer adaptive control strategies, such as similarity judgment [15, 24], knowledge selection [18, 25], and historical feedback [26, 27], have also been studied. There is a growing number of MTEAs proposed by researchers, yet there is no standardized programming language, code pattern, or software platform for the EMT source codes. This presents difficulties for beginners to enter the field of EMT and for researchers to make convenient experimental comparisons of the algorithms.
As EMT is booming, it has become urgent to provide a convenient and user-friendly comparison software platform for benchmarking the performance of MTEAs. Moreover, the available source code and platform are essential for exploring real-world applications of EMT. Open-source and user-friendly software platforms can greatly facilitate the development of a research field. As for EC, some popular and successful software platforms helped researchers advance the development of EA and evolutionary optimization. PlatEMO [28] is developed for evolutionary multi-objective optimization and benchmarking multi-objective EAs. It facilitates researchers to design multi-objective EAs to deal with various types of multi-objective optimization problems. EvoX [29] is a distributed GPU-accelerated library that helps researchers design parallel EAs to accelerate the solving of complex optimization problems and training for reinforcement learning. However, different from the traditional single-task EC field, solving MTO problems with MTEAs requires simultaneous evolution of multiple optimization tasks with inter-task solution space mapping and knowledge transfer. Additionally, the performance metrics of EMT are diverse, with both single-task metrics and multitask metrics. These special requirements are hard to implement in existing tools and platforms. Moreover, existing research work on EMT presents various code implementations of algorithms and problems. Particularly, EMT has been extended to the subfields of many-task optimization [30, 31], multi-objective multitask optimization [32, 33], competitive multitask optimization [26, 34], and constrained multi-task optimization [35]. The researchers cannot easily benchmark MTEAs on various types of MTO problems.
In order to address these issues, we introduce an open-source MATLAB platform, named MTO-Platform (MToP), for the development of the EMT field. The main contributions of MToP are summarized as follows:
  1. 1.
    MToP has an easy-to-use graphical user interface (GUI) including test, experiment, and data-process modules. These modules can assist researchers in realizing problem characteristics, performing comparative experiments, parallel execution, statistically analyzing results, plotting result figures, and processing experimental data. Moreover, MToP provides a modular implementation of algorithms, problems and metrics. On top of that, extensive public application programming interface (API) functions are provided, such as population initialization, function evaluation, evolutionary operator, and environmental selection.
  2. 2.
    An abundance of algorithms, problems, and metrics are implemented with public API and included in MToP. More than 30 MTEAs (see Table I) cover single-objective, multi-objective, constrained, many-task, and competitive EMT types. Additionally, to facilitate the comparison of MTEAs with popular traditional EAs, more than 30 single-task EAs (see Table II) of various types, such as genetic algorithm, differential evolution, swarm intelligence, and evolution strategy, are also modified and included in MToP. For benchmark test problems, more than 150 benchmark MTO problems and several real-world applications (see Table III) are incorporated. Finally, various types of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, are also implemented in MToP.
  3. 3.
    MToP is extensible to easily adding new code of algorithms, problems, and metrics. After following the coding patterns and implementing functionality based on the public API, new code can be recognized and performed. Since MToP is completely open-source, researchers can utilize existing algorithms, problems, and metrics as the basis for new ideas. Through the open-source website GitHub, MToP will be continuously updated and receive new code [^1] MToP has been continuously updated since Sep. 21, 2021, when it was uploaded to GitHub at https://github.com/intLyc/MTO-Platform..
TABLE I: Multitask evolutionary algorithms included in MToP (current version).
Algorithm Year Publication Objective Task Framework Constraint Description
MFEA [14] 2016 TEVC Single Multi Multifactorial Yes Multifactorial evolution-based GA for multitask optimization
LDA-MFEA [36] 2017 CEC Single Multi Multifactorial - Linearized domain adaptation in MFEA
MFDE&MFPSO [37] 2017 CEC Single Multi Multifactorial Yes Multifactorial evolution-based DE and PSO
IMEA [38] 2018 GECCO Single Multi Multi-population Yes Multi-population framework for multitask optimization through island model
G-MFEA [39] 2019 TEVC Single Multi Multifactorial - Generalized MFEA with decision variable translation and shuffling
EMEA [19] 2019 TCYB Single Multi Multi-population Yes Explicit knowledge transfer for multitask optimization via autoencoding
MFEA-GHS [40] 2019 ESWA Single Multi Multifactorial Yes Genetic Transform and hyper-rectangle search in MFEA
MFEA-DV [41] 2019 CEC Single Multi Multifactorial Yes Enhanced MFEA with cross-task search direction
SBO [31] 2019 AAAI Single/Multi Many Multi-population Yes Symbiosis in biocoenosis framework for many-task optimization
MFEA-II [15] 2020 TEVC Single Multi Multifactorial Yes MFEA with online transfer parameter estimation
SREMTO [42] 2020 TEVC Single Multi Multifactorial Yes Self-regulated multitask framework for varying relatedness among tasks
MFMP [17] 2020 INS Single Multi Multi-population Yes Multi-population-based adaptive DE for multitask optimization
TLTLA [43] 2020 FNS Single Multi Multifactorial Yes MFEA with two-level inter- and intra-task transfer learning
MaTDE [30] 2020 TETCI Single/Multi Many Multi-population Yes Many-task DE with adaptive archive-based knowledge transfer
MFEA-AKT [16] 2021 TCYB Single Multi Multifactorial - MFEA with adaptive knowledge transfer via multiple crossover operator
MTEA-AD [18] 2022 TEVC Single Multi Multi-population Yes MTEA with adaptive knowledge transfer via anomaly detection
DEORA [26] 2022 TEVC Single Multi Multi-population - Adaptive task selection for competitive multitask optimization
AT-MFEA [21] 2022 TCYB Single Multi Multifactorial Yes Affine transformation-enhanced domain adaptation for MFEA
MFEA-DGD [44] 2022 TCYB Single Multi Multifactorial - MFEA based on diffusion gradient descent
MTES [45] 2022 TCYB Single Multi Multi-population Yes Multitask OpenAI-ES via gradient-free evolution multitasking
BoKT [24] 2022 TEVC Single Many Multi-population - Bi-objective knowledge transfer framework for many-task optimization
EMaTO-MKT [25] 2022 TEVC Single/Multi Many Multi-population - Multi-source knowledge transfer via local distribution estimation
MTES-KG [46] 2023 TEVC Single Multi Multi-population Yes Multitask ES with knowledge-guided external sampling
BLKT-DE [47] 2023 TCYB Single Multi Multi-population Yes Multitask DE with block-level knowledge transfer
MTEA-SaO [48] 2023 INS Single/Multi Multi Multi-population Yes Adaptive solver multitask framework with implicit knowledge transfer
MTSRA [34] 2023 ESWA Single Multi Multi-population - Improved adaptive DE with competitive task selection
MO-MFEA [32] 2017 TCYB Multi Multi Multifactorial Yes Multiobjective MFEA with non-dominanted sort and crowding distance
MM-DE [49] 2018 CEC Multi Multi Multi-population Yes Fast memetic DE for multiobjective multitask optimization
AMT-NSGA-II [50] 2019 TCYB Multi Multi Multi-population Yes Curbing negative influences online for seamless transfer
MO-MFEA-II [51] 2021 TCYB Multi Multi Multifactorial Yes Cognizant Multitasking for parameter estimation in MO-MFEA
EMT-ET [27] 2021 TCYB Multi Multi Multi-population Yes Effective knowledge transfer approach via non-dominated sort
EMT-PD [33] 2021 SMCA Multi Multi Multifactorial Yes Two-stage adaptive knowledge transfer based on population distribution
MOMFEA-SADE [52] 2022 TCYB Multi Multi Multi-population Yes Subspace alignment and adaptive Differential Evolution
EMT-GS [23] 2023 TEVC Multi Multi Multifactorial Yes Generative adversarial networks for knowledge transfer
MTEA-D-DN [53] 2023 TEVC Multi Multi Multi-population - Neighborhood as a bridge for decomposition-based knowledge transfer
TABLE II: Single-task evolutionary algorithms for multitask optimization included in MToP (current version).
Algorithms Year Publication Objective Category Constraint Description
GA - - Single Genetic Algorithm Yes Genetic algorithm with SBX crossover and polynomial mutation
DE - - Single Differential Evolution Yes Differential evolution algorithm with DE/rand/1/bin operator
PSO [54] 1995 ICNN Single Swarm Intelligence Yes Particle swarm optimization with global and particle best update
CMA-ES [55] 2001 ECJ Single Evolution Strategy Yes Evolution strategy with derandomized covariance matrix adaptation
jDE [56] 2006 TEVC Single Differential Evolution Yes Self-adapting control parameters in DE
MTV-DE [57] 2007 EngOpt Single Constraint Handling Yes Multiple trial vectors in DE for handling constraints
JADE [58] 2009 TEVC Single Differential Evolution Yes Adaptive DE with fast convergence performance
ECHT [59] 2010 TEVC Single Constraint Handling Yes Ensemble of constraint handling techniques
xNES-as [60] 2012 GECCO Single Evolution Strategy Yes Natural ES with adaptation sampling
rank-DE [61] 2013 TCYB Single Differential Evolution Yes Ranking-based mutation operators for DE
SHADE [62] 2013 CEC Single Differential Evolution Yes Success-history based parameter adaptation for DE
R1-NES [63] 2013 GECCO Single Evolution Strategy Yes A Linear Time natural ES for non-separable functions
xNES&SNES [64] 2014 JMLR Single Evolution Strategy Yes ES with adaptive natural gradients to update distribution
L-SHADE [65] 2014 CEC Single Differential Evolution Yes Linear population size reduction for SHADE
CSO [66] 2015 TCYB Single Swarm Intelligence Yes Competitive swarm optimizer for large-scale optimization
FROFI [67] 2016 TCYB Single Constraint Handling Yes Incorporating objective function information into the feasibility rule
jSO [68] 2017 CEC Single Differential Evolution Yes Champion of CEC 2017 numerical optimization competition via improved DE
LSHADE44 [69] 2017 CEC Single Differential Evolution Yes L-SHADE with competing strategies applied to constrained optimization
CAL-SHADE [70] 2017 CEC Single Differential Evolution Yes Adaptive constraint handling technique for L-SHADE
OpenAI-ES [71] 2017 ArXiv Single Evolution Strategy Yes Parallelized ES with standard normal distribution gradients
C2oDE [72] 2019 SMCA Single Constraint Handling Yes Composite DE for constrained optimization
DES [73] 2020 TEVC Single Evolution Strategy Yes Matrix-free covariance matrix adaptation ES
CORCO [1] 2020 TEVC Single Constraint Handling Yes Utilizing the correlation between constraints and objective function
MPA [74] 2020 ESWA Single Swarm Intelligence Yes Marine predators algorithm for numerical optimization
EO [75] 2020 KBS Single Swarm Intelligence Yes Equilibrium optimizer for numerical optimization
NL-SHADE-RSP [76] 2021 CEC Single Differential Evolution - Champion of CEC 2021 numerical optimization competition via improved DE
AO [77] 2021 CIE Single Swarm Intelligence Yes Aquila optimizer for numerical optimization
DeCODE [78] 2021 SMCA Single Constraint Handling Yes Decomposition-based multiobjective approach for constrained optimization
VMCH [79] 2022 TEVC Single Constraint Handling Yes Voting-mechanism-based ensemble of constraint handling techniques
CEDE-DRL [80] 2023 SWEVO Single Constraint Handling Yes Deep reinforcement learning assisted co-evolutionary DE
SPEA2 [81] 2001 EUROGEN Multi Non-dominated Sort Yes Improving the strength Pareto approach for multiobjective optimization
NSGA-II [82] 2002 TEVC Multi Non-dominated Sort Yes Multiobjective GA with non-dominated sort and crowding distance
MO-CMA-ES [83] 2007 ECJ Multi Evolution Strategy - Covariance matrix adaptation for multiobjective optimization
MOEA/D [3] 2007 TEVC Multi Decomposition - Decomposition-based multiobjective evolutionary algorithm
MOEA/D-DE [84] 2009 TEVC Multi Decomposition - MOEA/D for complex Pareto sets in multiobjective optimization
NSGA-III [4] 2014 TEVC Multi Non-dominated Sort Yes Reference-point-based nondominated sorting for multiobjective optimization
LMOCSO [85] 2020 TCYB Multi Swarm Intelligence Yes Efficient large-scale multiobjective optimization based on CSO
CCMO [2] 2021 TEVC Multi Constraint Handling Yes Coevolutionary framework for constrained multiobjective optimization
CMOCSO [86] 2022 TEVC Multi Swarm Intelligence Yes Competitive and cooperative swarm optimizer for multiobjective optimization
TABLE III: Multitask optimization problems included in MToP (current version).
Problem Case Task Dimension Objective Special Description
CEC17-MTSO [87] 9 2 50 Single - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MaTSO [88] 6 2-50 50 Single - Single-objective many-task optimization problems
WCCI20-MTSO [89] 10 2 50 Single - Complex single-objective multitask optimization problems
WCCI20-MaTSO [89] 10 2-50 50 Single - Complex single-objective many-task optimization problems
CMT [35] 9 2 Any Single Constrained Large/partial/no feasible regions interaction and different inter-task similarity
C2TOP&C4TOP [26] 51 2/4 50 Single Competitive Different optimal objective values among competitive tasks
C-CPLX [34] 20 2 50 Single Competitive Complex competitive multitask optimization problems
CEC17-MTMO [90] 9 2 50 Multi - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MTMO [88] 10 2 50 Multi - Complex multi-objective multitask optimization problems
CEC19-MaTMO [88] 10 2-50 50 Multi - Multi-objective many-task optimization problems
WCCI20-MaTMO [89] 10 2-50 50 Multi - Complex multi-objective many-task optimization problems
CEC21-MTMO [91] 10 2 50 Multi - More complex multi-objective multi optimization problems
PKACP [24] 1 Any Any Single - Planar kinematic arm control problem as many-task optimization problem
SCP [26] 1 Any 60-120 Single Competitive Sensor coverage problem as competitive multitask optimization problem
PEPVM [48] 1 3 7 Single - Parameter extraction of photovoltaic models as multitask optimization problem
OPF [46] 5 2 24/33 Single Constrained Optimal power flow as constrained multitask optimization problem
SOPM 2 3 30 Multi Constrained Synchronous optimal pulse-width modulation as constrained multiobjective multitask problem
MO-OPF 4 2 24/33 Multi Constrained Multiobjective optimal power flow as constrained multiobjective multitask problem
MO-SCP 1 Any 60-120 Multi Competitive Multiobjective sensor coverage problem as competitive multiobjective multitask problem
TABLE IV: Metrics included in MToP (current version).
Metric Objective Description
Obj Single Objective value for each task
Obj (MTS) Single Multitask score of Obj for all tasks
Obj (AV) Single Average Obj for all tasks
Obj (CMT) Single Competitive multitask Obj for all tasks
Obj (NBR) Single Number of best Obj result for all tasks
CV Single Constraint violation for each task
FR Single Feasible rate for each task
HV Multi Hypervolume for each task
HV (MTS) Multi Multitask score of HV for all tasks
HV (CMT) Multi Competitive multitask HV for all tasks
IGD Multi Inverted generational distance for each task
IGD+ Multi Improved IGD plus for each task
IGD+ (MTS) Multi Multitask score of IGD+ for all tasks
IGD+ (CMT) Multi Competitive multitask IGD+ for all tasks
Run Time All Algorithm running time for all tasks
The rest of this paper is organized as follows. Section II elaborates on the architecture of MToP. Section III describes approaches for performing experiments with MToP. Section IV introduces how to extend contents of MToP. Finally, the conclusion and future work are given in Section V.
II Architecture of MToP
II-A Functional Modules
II-A1 Test Module
(a) Function landscapes
(b) Feasible regions
(c) Pareto front
(d) Metric convergence
Figure 1: Examples of graphic display in the test module of MToP.
The test module is designed to facilitate researchers in analyzing the characteristics of MTO problems and algorithms. The test module of MToP provides functions to run the selected algorithm on the selected problem. It provides visualization functions for both single-objective and multi-objective MTO problems. For single-objective MTO problems, the landscapes of the objective function with different tasks in the one-dimensional unified search space can be displayed as in Fig. 1 (a). For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be shown as in Fig. 1 (b). For multi-objective MTO problems, the Pareto front figures can be displayed for multiple tasks as in Fig. 1 (c). The metric convergence behavior can be drawn after running algorithms on problems as in Fig. 1 (d).
Figure 2: The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
The workflow of API calls for the execution of an algorithm on a problem is displayed in Fig. 2. The workflow starts with the Start_Button event in MTO_GUI.m, which calls the parameter setting API Set_Parameter in the selected algorithm MFEA.m. Following this, it enters the Run function of the algorithm, which at the beginning calls the Initialization API of the algorithm base class Algorithm.m and later loops through the function evaluation API Evaluation. The Initialization gets the parameter settings of the problem CEC17_MTP1.m to be solved via the problem base class Problem.m. Similarly, the Evaluation gets the specific optimization task object through the problem base class. After the algorithm has run, GUI then displays the resultant performance metrics after calling the metrics function Obj.m via Call_Metric API. All functions, including the subsequent experiment module, in MToP that execute algorithms are based on this procedure. Both algorithm and problem parameters can be modified directly through the GUI of MToP.
II-A2 Experiment Module
TABLE V: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
GA MFEA MTEA-AD
CI-HS-T1 9.0426e-01 (6.16e-02) 8.4445e-01 (5.51e-02) + 9.3438e-01 (5.03e-02) =
CI-HS-T2 4.2400e+02 (4.90e+01) 2.8119e+02 (5.09e+01) + 3.1099e+02 (4.11e+01) +
CI-MS-T1 5.4562e+00 (1.52e+00) 5.8307e+00 (8.23e-01) - 4.5085e+00 (2.69e-01) +
CI-MS-T2 4.2474e+02 (6.45e+01) 3.1448e+02 (4.41e+01) + 3.1660e+02 (4.39e+01) +
CI-LS-T1 2.1167e+01 (1.82e-01) 2.0234e+01 (8.35e-02) + 2.1086e+01 (2.82e-01) =
CI-LS-T2 4.1476e+03 (6.34e+02) 4.3638e+03 (4.88e+02) = 4.5958e+03 (6.64e+02) -
PI-HS-T1 4.2186e+02 (5.27e+01) 6.6221e+02 (1.11e+02) - 4.4662e+02 (4.13e+01) -
PI-HS-T2 8.6994e+01 (1.93e+01) 8.9970e+01 (1.67e+01) = 1.2177e+02 (2.88e+01) -
PI-MS-T1 5.4736e+00 (1.42e+00) 4.9276e+00 (4.76e-01) = 4.7264e+00 (3.50e-01) +
PI-MS-T2 2.4152e+04 (1.00e+04) 1.0488e+04 (4.77e+03) + 9.1688e+03 (4.69e+03) +
PI-LS-T1 5.6540e+00 (1.88e+00) 1.9583e+01 (2.71e+00) - 5.3895e+00 (9.11e-01) =
PI-LS-T2 1.4062e+01 (2.48e+00) 1.9371e+01 (3.26e+00) - 5.1076e+00 (1.44e+00) +
NI-HS-T1 2.9452e+04 (1.49e+04) 9.7165e+03 (4.93e+03) + 1.3459e+04 (4.77e+03) +
NI-HS-T2 4.5183e+02 (6.26e+01) 3.7757e+02 (7.84e+01) + 3.8559e+02 (6.26e+01) +
NI-MS-T1 9.0462e-01 (6.19e-02) 9.3064e-01 (5.16e-02) = 9.8876e-01 (4.48e-02) -
NI-MS-T2 3.8662e+01 (3.97e+00) 2.7501e+01 (3.26e+00) + 2.4831e+01 (3.76e+00) +
NI-LS-T1 4.2895e+02 (5.41e+01) 7.4411e+02 (1.30e+02) - 4.6340e+02 (4.97e+01) -
NI-LS-T2 4.3308e+03 (5.19e+02) 4.5452e+03 (4.90e+02) = 4.7449e+03 (6.45e+02) -
+ / - / = Base 8 / 5 / 5 9 / 6 / 3
(a) Convergence
(b) Pareto front
Figure 3: schematic plotting by the experiment module of MToP.
The experiment module provides functions for performing comprehensive experiments and analyzing results. Users can execute multiple independent runs of multiple algorithms to solve multiple MTO problems. Data such as experimental settings, objective values, and decision variables of the execution are recorded and stored after running is completed. Based on these data, metric results, significance tests, and schematic plotting would be able to be continued. Metrics functions running on the data are divided into single-task and multitask metrics. Selecting single-task metrics displays the results of each task of the MTO problem on the table in MToP, and selecting multitask metrics displays the results of each MTO problem. The results of the significance tests are also shown in the table along with the metric results. All table data can be exported in tex, csv, and xlsx types as shown in Table V. Convergence plots are displayed based on the selected table data to analyze the comparison of the convergence behavior of different algorithms as in Fig. 3 (a). Additionally, for the multi-objective MTO algorithms and problems, the Pareto front can plotted as in Fig. 3 (b).
II-A3 Data Process Module
TABLE VI: Properties of experimental data.
Property Description
Reps Repetitions number of independent runs
Algorithms Algorithms data contains names and parameter settings
Problems Problems data contains names and parameter settings
Results Total results data contains Obj, CV, and Dec
RunTimes Running time data of algorithms on problems
Metrics Calculated metric results data
Each unique execution in the experiment module is treated as a data object and saved to a data file MTOData.mat. The settings and intermediate results of this experiment are stored in data properties as shown in Table VI. The saved data can subsequently be reloaded by the experiment module of MToP. With the data process module, the data can be merged or split which allows data to be reused and customized. The processing of data is grouped into three types, which are independent runs, algorithms and problems. Having several sets of data with fewer independent runs, the data can then be merged to become new data with a larger number of independent runs. In cases where existing data needs to be augmented with new algorithms for comparison, with the data process module there is no need to re-execute all the comparative experiments. It is only necessary to run the new algorithms and merge the resulting data with the original data. Similarly, if not all algorithms in the data are needed, it can be split and re-merged. As with the data processing for algorithms, the data processing for problems is applied in cases where new problems need to be added or deleted.
II-B File Structure
Figure 4: File structure of MToP.
In the root directory MTO/ of MToP, there is a running script file mto.m and four folders, Algorithms/, Problems/, Metrics/, and GUI/. The structure of these folders is shown in Fig. 4. The algorithms folder Algorithms/ holds various categories folder, utility functions, and the algorithm base class Algorithm.m. The specific algorithm files, e.g., MFEA.m, MO_MFEA.m, GA.m, are stored under their corresponding classification folder. The utility functions folder Algorithms/Utils/ provides various public utility functions of generation operators, environmental selection approaches, constraint handling techniques, etc. The problems folder Problems/ contains various categories folder, base function folder, and the problem base class Problem.m. The specific problem files, e.g., WCCI20_MTSO1.m, CMT3.m, CEC19_MaTMO6.m, are stored under their corresponding classification folder. In particular, files for real-world application problems are placed in Real-World Application/ folder. The Problems/Base/ folder includes the public base functions for problems. The third folder Metrics/ in MToP contains all the result evaluation metrics and is with two subfolders, Metrics/Single-objective/ and Metrics/Multi-objective for single-objective and multi-objective optimization respectively. Specific metric files are Obj.m, Obj_MTS.m, IGD.m, etc. The last folder is the GUI/ which contains all the files used by the GUI of MToP. Among them, MTO_GUI.m is the main file of the GUI, and MTO_CMD.m provides the function of executing experiments by the command line.
II-C Code Pattern
II-C1 Algorithm
TABLE VII: Properties and methods of algorithm base class.
Property or method Description
FE Number of fitness function evaluations
Gen Number of evolutionary generations
Best Best individual found for single-objective optimization
Result Result data contains Obj, CV, and Dec
getParameter() Get customized parameters for algorithm object
setParameter() Set customized parameters for algorithm object
notTerminated() Determine whether to terminate and update result data
Evaluation() Fitness function evaluation
run() Executing algorithm
1    label={list:algo},
2    caption={Algorithm implementation example.},
3    captionpos=b,
4    float=htbp
5]
6classdef Algo_Example < Algorithm
7% <Multi-task> <Single-objective> <None/Constrained>
8
9properties % set parameter
10  para_example = 0.5;
11end
12
13methods
14  function run(Algo, Prob)
15    % initialize populations for all tasks
16    pops = Initialization(Algo, Prob, Individual);
17
18    while Algo.notTerminated(Prob, pops) % main loop
19      for t = 1:Prob.T % for each task
20        % generate new offspring
21        off = Generation(pops{t});
22        % evaluate the fitness of offspring
23        off = Algo.Evaluation(off, Prob, t);
24        % environmental selection
25        pops{t} = Selection(pops{t}, off);
26      end
27    end
28  end
29end
All algorithms in MToP inherit from the algorithm base class Algorithm.m. The properties and methods of the algorithm base class are shown in Table VII. The base class provides all the functions that interact with the GUI, so the implemented algorithms only need to consider the evolutionary workflow itself. A simple example of algorithm implementation is shown in Listing LABEL:list:algo. The second line here is labels for the algorithms, inspired by PlatEMO [28], for the algorithm classification displayed in the GUI. The function run() of the class is for the execution of the algorithm, with its first argument Algo denoting the object itself and the second argument Prob denoting the problem object to be solved. At the beginning of the run() function, the population initialization is performed with the Initialization() function. MToP provides two population initialization functions, the multi-population method and the multifactorial method, which produce either multiple populations or single populations with skill factors respectively. Here the pops produced in line 11 are of cell data type containing multiple populations. Each individual in the pops is Individual object and contains properties including objective value Obj, constraint violation CV, and decision variables Dec. Then, the main loop is entered with the function notTerminated() implemented in the algorithm base class. The function notTerminated() provides data updates and generation counting methods. In the main loop, the functions of offspring generation Generation(), offspring evaluation Evaluation(), and environmental selection Selection() are performed separately for each task. For algorithms with the multifactorial framework, these functions are executed simultaneously for all tasks. The function Generation() acts on the decision variables Dec of offspring individuals and is implemented by specific algorithms. The evaluation function Evaluation() is implemented in the algorithm base class, which automatically calls the problem object Prob for offspring fitness evaluation, i.e., calculating the objective value Obj and constraint violation CV for all individuals. It also updates the best solution automatically for single-objective optimization. At the end of each loop, through the environmental selection function Selection(), the new populations pops are updated. MToP provides some universal Selection() functions, and this function can also be reimplemented by specific algorithms. After that, the code goes to the next loop and calls notTerminated() to record changes. The algorithm code pattern in MToP is adapted to all EAs for solving MTO problems.
II-C2 Problem
TABLE VIII: Properties and methods of problem base class.
Property or method Description
T Number of optimization tasks
N Population size for each task
M Number of objectives for all tasks
D Number of decision variable dimensions for all tasks
Fnc Fitness function for all tasks
Lb Lower bound of decision variables for all tasks
Ub Upper bound of decision variables for all tasks
maxFE Maximum fitness function evaluations
getRunParameter() Get public parameters for problem object
getParameter() Get customized parameters for problem object
setParameter() Set customized parameters for problem object
setTasks() Set optimization tasks
getOptimum() Get optimum solutions for multi-objective optimization
1    label={list:prob},
2    caption={Problem implementation example.},
3    captionpos=b,
4    float=htbp
5]
6classdef Prob_Example < Problem
7% <Multi-task> <Multi-objective> <None>
8
9methods
10function Prob = Prob_Example(varargin)
11    Prob = Prob@Problem(varargin);
12    % set default maximum function evaluations
13    Prob.maxFE = 1000 * 100;
14  end
15
16  function setTasks(Prob)
17    Prob.T = 2;
18    % task 1
19    Prob.D(1) = 10; % variable dimensions
20    Prob.M(1) = 2; % objective number
21    Prob.Fnc{1} = @func1 % fitness function
22    Prob.Lb{1} = zeros(1, 10); % lower bound
23    Prob.Ub{1} = ones(1, 10); % upper bound
24    % task 2
25    Prob.D(2) = 20; % variable dimensions
26    Prob.M(2) = 3; % objective dimensions
27    Prob.Fnc{2} = @func2 % objective function
28    Prob.Lb{2} = [0,-ones(1, 19)]; % lower bound
29    Prob.Ub{2} = [1,ones(1, 19)]; % upper bound
30  end
31
32  function optimum = getOptimum(Prob)
33    % return optimum points for each task
34    optimum{1} = optimum_matrix1;
35    optimum{2} = optimum_matrix2;
36  end
37end
All problems in MToP inherit from the problem base class Problem.m. The properties and methods of the Problem.m are shown in Table VIII. A simple example of problem implementation is presented in Listing LABEL:list:prob. The maximum number of function evaluations maxFE for the problem can be defined in the constructor of the class. The function setTasks() is for setting the properties and evaluation function for each optimization task. In this function, the number of tasks T of the problem is first defined followed by a detailed per-task setup. After that, decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub need to be set separately for each task. M and D for multiple tasks of the problem are denoted as vector type. Fnc, Lb and Ub are denoted as cell type. Note that the function Fnc for each task takes inputs of the decision variable Dec and returns objective values Obj and the constraint violation value CV. Additionally, for multi-objective optimization, the external function getOptimum(), which obtains optimal solutions (i.e., reference points), needs to be defined. With these problem code patterns, all problems that are completely defined can be solved using the corresponding type of algorithm in MToP.
II-C3 Metric
1    label={list:metric},
2    caption={Metric implementation example.},
3    captionpos=b,
4    float=htbp
5]
6function result = Metric_Example(MTOData)
7% <Metric> <Multi-objective>
8
9result.Metric = ’Min’;
10
11% Data for shown in the GUI table
12result.RowName = {MTOData.Problems.Name};
13result.ColumnName = {MTOData.Algorithms.Name};
14result.TableData = CalculateTableData(MTOData);
15
16% Data for shown in the GUI convergence plot
17result.ConvergeData = CalculateConvergeData(MTOData);
18
19% Data for shown in the GUI Pareto plot
20result.ParetoData = CalculateParetoData(MTOData);
21end
In contrast to algorithm and problem class implementation, metric codes in MToP are recognized as the function type. A simple example of metric implementation is shown in Listing LABEL:list:metric. The input parameter of the function is MTOData generated by the experimental execution. The function returns result, which contains properties of Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData. The property Metric can be either ‘Min’ or ‘Max’, indicating that a smaller or larger metric value is better. RowName, ColumnName, and TableData are used to return the metric results for the GUI table. ConvergeData and ParetoData are used to return metric convergence results for convergence plots and non-dominated solutions for Pareto plots respectively. Note that specific metric calculation functions are not covered here for the simplicity of this example.
III Performing Experiments with MToP
(a) Test module
(b) Experiment module
Figure 5: Graphical user interface of MToP.
Running MToP requires first executing the script file mto.m in the root directory. After that, the MToP GUI is shown as presented in Fig. 5.
III-A Testing Algorithms and Problems
As shown in Fig. 5 (a), the test module is divided into left, center, and right parts. The left part is for the selection and setup of algorithms and problems. After selecting the corresponding labels of Task, Objective, and Special drop-down buttons, the Algorithm and Problem drop-down buttons update their contents. Selecting specific algorithms and problems displays the parameters they contain in the edit box, which can be also clicked on and set. After parameters are updated, new parameter values are fed back into the algorithm and problem objects through the setParameter() method. The parameters of the algorithms generally contain evolutionary operator parameters and their customized parameters. The problem parameters generally contain population size N, maximum function evaluations maxFE, and their customized parameters. The center part of the MToP GUI can execute the test experiment and display the graphic results. After executing the procedure of the algorithm on the problem, the metric results of the problem are displayed in the right part of the test module.
III-B Executing Comparative Experiments
The GUI of the experiment module is shown in Fig. 5 (b). It is divided into a two-column experimental setup part on the left and a results display part on the right. The first column contains the basic settings to perform experiments. The text box No. of Runs indicates the number of independent repeat runs for each algorithm on each problem. The text box No. of Results denotes the number of results saved by the evolutionary process of each algorithm, which is uniformly sampled according to the number of generations. The Parallel button controls whether experiments are executed in parallel processing. Parallel execution invokes multiple threads to run simultaneously multiple independent runs. The drop-down buttons Task, Objective, and Special serve the same purpose as the test module, filtering algorithms and problems based on their pre-defined labels. After selecting the algorithm and problem in the two boxes on the bottom side of the first column, it can be added to the preparation area of the second column by clicking the Add button. The names and parameters of the added algorithms and problems can be edited by clicking on them. After these are ready, click on the green Start button and MToP starts executing the experiment. It is possible to click on the buttons Pause and Stop to pause or stop the experiment during the execution.
When the experiment has been completed, the content of the table is available in the results area on the right part of the experiment module. The original data of the experiment can be read and saved by Load Data and Save Data buttons. In the right part, The buttons in the first function row operate on the experimental data. The first button Save Table in the line can export the results of the table to xlsx, csv, and tex files. The second and third Converge buttons control convergence y-lim types and draw convergence plots based on the results selected in the table. The fourth button Pareto draws Pareto front and non-dominated solutions based on the selected results. The fifth edit box controls the type of result formatting of the table. The drop-down box right of the edit box is for selecting performance metrics. The drop-down boxes following the metrics are the result display type, statistical test method, base algorithm, and result highlighting type, respectively. After the statistical test method is selected, the calculated statistical results, denoted by “+/-/=” followed by each metric result, can be also displayed in the table. MToP provides two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests. The base algorithm drop-down button controls the base algorithm in the statistical test. The highlight button highlights the best and worst algorithm’s metrics results for each problem or task.
III-C Executing via Alternative Command Line
In addition to GUI execution, MToP provides alternative options for running experiments via the command line. After inputting the algorithm list, problem list and other parameters into the script file mto.m, the experiments can be run directly in the command line and the data MTOData.mat are returned after running completed. Similarly, inputting the data into the metrics function yields the calculated metric results.
IV Extending MToP
New algorithms, problems and performance metrics can be easily added in MToP. Users can create and implement code files following the code pattern described in Section II-C.
IV-A Adding New Algorithms
The first step in adding an algorithm to MToP is to create a MATLAB class file that inherits the base class Algorithm.m in the Algorithms/ folder. The algorithmic framework can then be implemented following the code pattern as in Listing LABEL:list:algo. Users can use the population initialization function Initialization(), offspring generation function Generation(), and environmental selection function Selection() available in MToP, or of course reimplement these functions according to their requirements. In addition, users who want to create customized parameters that can be clicked and modified in the MToP GUI require an additional implementation of getParameter() and setParameter() API in the new algorithm class file.
IV-B Adding New Problems
Similar to adding algorithms, adding problems in MToP requires the creation of a specific problem class file that inherits the base class Problem.m in the Problems/ folder. The suggested maximum function evaluations maxFE need to be predefined in the constructor of the specific problem. After that, refer to Listing LABEL:list:prob to implement the configuration of each optimization task in setTasks() function. When optimization tasks include multi-objective optimization problems, it is necessary to additionally implement the getOptimum() function to return the optimum solution set or reference points. Finally, for setting customized parameters that can be modified in the GUI, it is also required to implement getParameter() and setParameter() functions.
IV-C Adding New Performance Metrics
Users can add new performance metrics by creating a new metrics function file in the Metrics/ folder. Follow the pattern as in Listing LABEL:list:metric to receive the data structure object MTOData and return the calculated result object. The metric results displayed in the GUI table need to return the RowName, ColumnName, and TableData fields. The RowName for combined multitask metrics is typically the problem name, while metrics on individual tasks need to return the name of each task. The ColumnName is typically the algorithm name. The calculation function of TableData needs to be implemented for the specific performance metric. To implement the function to plot the convergence of the metric, the ConvergeData field is also required. Additionally, for multi-objective optimization, the metric function should also return non-dominated solutions in the ParetoData field.
V Conclusion and Future Work
This paper introduced an open-source MATLAB platform MToP for EMT. MToP provides a user-friendly GUI, enriched algorithms and problems, and convenient code patterns. The current version of MToP includes more than 30 MTEAs, more than 30 single-task EAs (that can handle MTO problems), more than 150 MTO benchmark problems, and several real-world applications of EMT. The functional modules, file structure and code pattern of MToP are described in detail. Moreover, the procedures and extensions for conducting experiments using MToP are also presented.
The current MToP, while carefully reimplementing, modifying, and testing the included algorithms and problems, might still have some reproducing errors and bugs. We have been actively looking for and fixing the issues and continue to update MToP. Fortunately, MToP is uploaded on the open-source website GitHub, where followers and researchers can test and give feedback on issues. After the upload, we also received and updated many codes from researchers.
In the future, we will continue to develop and update MToP with inspiration and contributions from the academic community. We hope that MToP can become a convenient tool to help researchers advance the field of EMT and consequently advance the evolutionary computation field.
Acknowledgement
The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO [28], for which the authors are also thankful.