### v2

# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

Abstract
Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past years. It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques. Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 10 performance metrics. Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 40 popular single-task evolutionary algorithms to address MTO problems. MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization. More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
Index Terms:
Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm.
I Introduction
Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency. Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and user-friendly nature. Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6]. In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking [7]. A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ optimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
$\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,F_{K}(\bm{x}_{K})\Big{]},\\ \textit{s.t. }\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$ (1)
where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective optimization task, the $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task. EMT has found successful applications in various domains, including engineering scheduling [5, 8], nonlinear equation systems [9], feature selection in machine learning [10], anomaly detection [11], point cloud registration [12, 13], and reinforcement learning [14].
To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques [15, 16]. Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources [17, 18]. The first attempt of EMT can be traced back to the multifactorial EA [19], which introduced an implicit knowledge representation via random mating among optimization tasks. Subsequently, numerous EAs tailored for MTO have emerged. These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks [20, 21], or a multi-population framework allocating separate populations for each task [22, 23]. Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space [19, 21], autoencoding [24, 25], affine transformation [26, 27, 28], and adversarial generative models [29] have been proposed in MTEAs. Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment [20, 30, 31], knowledge selection [23, 32], and historical feedback [33, 34, 35], have also been investigated. Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes. This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs. Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field. Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field. In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization. For instance, PlatEMO [36] is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems. EvoX [37] is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training. Evo-Panel [38] created a comprehensive dynamic visualization tool to illustrate the detailed procedures of different optimization algorithms in solving numerical benchmark functions. EDOLAB [39] is an optimization tool specifically for dynamic optimization.
However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks. Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics. These distinctive requirements pose unsolvable challenges for MTEA implementation in existing tools and platforms. Moreover, EMT has been extended to the subfields of many-task optimization [40, 41], multi-objective multitask optimization [42, 43], competitive multitask optimization [33, 44], and constrained multi-task optimization [45]. Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. The main contributions of MToP are summarized as follows:
  1. 1.
    MToP features a user-friendly graphical user interface (GUI) comprising test, experiment, and data-process modules. These modules facilitate researchers in understanding problem characteristics, conducting comparative experiments, solving problems in parallel, statistically analyzing results, plotting result figures, and managing experimental data. Moreover, MToP offers a modular implementation of algorithms, problems and performance metrics. On top of these, it provides an extensive public application programming interface (API) with template functions for population initialization, function evaluation, evolutionary operators, and environmental selection.
  2. 2.
    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API. Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types (see Table IV). Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types (see Table V). In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications (see Table VII). Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities (see Table VIII).
  3. 3.
    MToP is designed to be easily extended, allowing for the seamless addition of new algorithms, problems, and metrics. By adhering to established coding patterns and implementing functionality based on the public API, new code can be seamlessly integrated and utilized within MToP. Given its status as a completely open-source project, researchers have the opportunity to leverage existing algorithms, problems, and metrics as the foundation for novel ideas. Through the collaborative platform GitHub, MToP undergoes continuous updates and enhancements, ensuring that it remains at the forefront of EMT research [^1] MToP has been continuously updated since Sep. 21, 2021, when it was uploaded to GitHub at https://github.com/intLyc/MTO-Platform..
The rest of this paper is organized as follows. Section II elaborates on the architecture of MToP. Section III describes approaches for performing experiments with MToP. Section IV introduces how to extend the contents of MToP. Finally, the conclusion and future work are given in Section V.
II Architecture of MToP
II-A Functional Modules
II-A1 Test Module
(a) Function landscapes (1D)
(b) Function landscapes (2D)
(c) Feasible regions
(d) Pareto front
(e) Metric convergence
Figure 1: Examples of graphic display in the test module of MToP.
The test module is designed to assist researchers in analyzing the characteristics of MTO problems and algorithms. It facilitates running the selected algorithm on the chosen problem within MToP. The module includes visualization functions for both single-objective and multi-objective MTO problems. In the case of single-objective MTO problems, the landscapes of the objective function with different tasks in the one- and two-dimensional unified search space can be depicted, as shown in Fig. 1 (a)-(b). For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be visualized, as illustrated in Fig. 1 (c). Regarding multi-objective MTO problems, the Pareto front figures can be generated for multiple tasks, as depicted in Fig. 1 (d). Furthermore, the convergence behavior of metrics can be plotted after executing algorithms on problems, as exemplified in Fig. 1 (e).
(a) Gen=1
(b) Gen=100
(c) Gen=200
(d) Gen=500
(e) Gen=700
(f) Gen=1000
Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space with the number of generations.
(a) Gen=1
(b) Gen=100
(c) Gen=200
(d) Gen=500
(e) Gen=700
(f) Gen=1000
Figure 3: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the decision space with the number of generations.
In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior. Figures 2 and 3 illustrate an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space. Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space. As evolution progresses through the first and middle phases (Gen=100,200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity. Subsequently, in the middle and late stages (Gen=500,700), the population gradually emphasizes diversity. By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task. The schematics indicate that the population for the second task occupies a similar position in the decision space as the first task. However, these two tasks do not precisely overlap in the decision space. This observation suggests the presence of negative knowledge transfer, which could be contributing to this discrepancy. In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
Figure 4: The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
The workflow of API calls for executing an algorithm on a problem is depicted in Fig. 4. The process initiates with the Start_Button event in MTO_GUI.m, triggering the parameter setting API Set_Parameter within the selected algorithm MFEA.m. Subsequently, it proceeds to the Run function of the algorithm, which initially invokes the Initialization API of the algorithm base class Algorithm.m, and then iterates through the function evaluation API Evaluation. During the initialization phase, the parameter settings of the problem CEC17_MTP1.m to be solved are obtained via the problem base class Problem.m. Similarly, the Evaluation function retrieves the specific optimization task object through the problem base class. Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function Obj.m via the Call_Metric API.
All functions within MToP that execute algorithms, including subsequent experiment modules, adhere to this procedure. Additionally, both algorithm and problem parameters can be directly modified through the GUI of MToP.
II-A2 Experiment Module
TABLE I: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
GA MFEA MTEA-AD
CI-HS-T1 9.0426e-01 (6.16e-02) 8.4445e-01 (5.51e-02) + 9.3438e-01 (5.03e-02) =
CI-HS-T2 4.2400e+02 (4.90e+01) 2.8119e+02 (5.09e+01) + 3.1099e+02 (4.11e+01) +
CI-MS-T1 5.4562e+00 (1.52e+00) 5.8307e+00 (8.23e-01) - 4.5085e+00 (2.69e-01) +
CI-MS-T2 4.2474e+02 (6.45e+01) 3.1448e+02 (4.41e+01) + 3.1660e+02 (4.39e+01) +
CI-LS-T1 2.1167e+01 (1.82e-01) 2.0234e+01 (8.35e-02) + 2.1086e+01 (2.82e-01) =
CI-LS-T2 4.1476e+03 (6.34e+02) 4.3638e+03 (4.88e+02) = 4.5958e+03 (6.64e+02) -
PI-HS-T1 4.2186e+02 (5.27e+01) 6.6221e+02 (1.11e+02) - 4.4662e+02 (4.13e+01) -
PI-HS-T2 8.6994e+01 (1.93e+01) 8.9970e+01 (1.67e+01) = 1.2177e+02 (2.88e+01) -
PI-MS-T1 5.4736e+00 (1.42e+00) 4.9276e+00 (4.76e-01) = 4.7264e+00 (3.50e-01) +
PI-MS-T2 2.4152e+04 (1.00e+04) 1.0488e+04 (4.77e+03) + 9.1688e+03 (4.69e+03) +
PI-LS-T1 5.6540e+00 (1.88e+00) 1.9583e+01 (2.71e+00) - 5.3895e+00 (9.11e-01) =
PI-LS-T2 1.4062e+01 (2.48e+00) 1.9371e+01 (3.26e+00) - 5.1076e+00 (1.44e+00) +
NI-HS-T1 2.9452e+04 (1.49e+04) 9.7165e+03 (4.93e+03) + 1.3459e+04 (4.77e+03) +
NI-HS-T2 4.5183e+02 (6.26e+01) 3.7757e+02 (7.84e+01) + 3.8559e+02 (6.26e+01) +
NI-MS-T1 9.0462e-01 (6.19e-02) 9.3064e-01 (5.16e-02) = 9.8876e-01 (4.48e-02) -
NI-MS-T2 3.8662e+01 (3.97e+00) 2.7501e+01 (3.26e+00) + 2.4831e+01 (3.76e+00) +
NI-LS-T1 4.2895e+02 (5.41e+01) 7.4411e+02 (1.30e+02) - 4.6340e+02 (4.97e+01) -
NI-LS-T2 4.3308e+03 (5.19e+02) 4.5452e+03 (4.90e+02) = 4.7449e+03 (6.45e+02) -
+ / - / = Base 8 / 5 / 5 9 / 6 / 3
(a) Convergence
(b) Pareto front
Figure 5: schematic plotting by the experiment module of MToP.
The experiment module offers functionalities for conducting comprehensive experiments and analyzing results. Users can execute multiple independent runs of various algorithms to tackle multiple MTO problems. Upon completion of the runs, data including experimental settings, objective values, and decision variables are recorded and stored. These data serve as the basis for generating metric results, conducting significance tests, and creating schematic plots.
Metrics functions operating on the data are categorized into single-task and multitask metrics. The selection of single-task metrics presents the results of each task of the MTO problem in tabular format within MToP while choosing to multitask metrics displays the results of each MTO problem. Significance test results are also included in the table alongside the metric results. All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table I.
Convergence plots are generated based on the selected table data to facilitate the comparison of convergence behaviors among different algorithms, as depicted in Fig. 5 (a). The convergence plot function can be applied to both single- and multi-objective problems. Furthermore, for multi-objective MTO algorithms and problems, the Pareto front can be plotted, as illustrated in Fig. 5 (b).
II-A3 Data Process Module
TABLE II: Properties of experimental data.
Property Description
Reps Repetitions number of independent runs
Algorithms Algorithms data contains names and parameter settings
Problems Problems data contains names and parameter settings
Results Total results data contains Obj, CV, and Dec
RunTimes Running time data of algorithms on problems
Metrics Calculated metric results data
Each unique execution within the experiment module is considered a data object and is saved to a data file named MTOData.mat. The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table II. This saved data can subsequently be reloaded by the experiment module of MToP.
With the data process module, the data can be merged or split, enabling data reuse and customization. The processing of data is categorized into three types: independent runs, algorithms, and problems. If several sets of data with fewer independent runs are available, the data can be merged to create new data with a larger number of independent runs. Similarly, in cases where existing data needs augmentation with new algorithms for comparison, it is unnecessary to re-execute all comparative experiments. Instead, only the new algorithms need to be executed, and the resulting data can be merged with the original data. Likewise, if not all algorithms within the data are required, the data can be split and re-merged accordingly. The data processing for problems follows a similar approach, wherein new problems can be added or deleted as needed.
II-B File Structure
Figure 6: File structure of MToP.
In the root directory MTO/ of MToP, you’ll find a script file named mto.m and four folders: Algorithms/, Problems/, Metrics/, and GUI/. The structure of these folders is depicted in Fig. 6.
The Algorithms/ folder contains various categories of algorithms, utility functions, and the algorithm base class Algorithm.m. Specific algorithm files such as MFEA.m, MO_MFEA.m, and GA.m are stored within their respective classification folders. The Algorithms/Utils/ folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
The Problems/ folder houses various categories of problems, the base function folder, and the problem base class Problem.m. Specific problem files like WCCI20_MTSO1.m, CMT3.m, and CEC19_MaTMO6.m are stored under their corresponding classification folders. Real-world application problem files are located in the Real-World Application/ folder, while the Problems/Base/ folder contains public base functions for problems.
The Metrics/ folder contains all result evaluation metrics, organized into subfolders for single-objective and multi-objective optimization (Metrics/Single-objective/ and Metrics/Multi-objective). Specific metric files include Obj.m, Obj_MTS.m, and IGD.m, among others.
Lastly, the GUI/ folder houses all files used by the GUI of MToP. Among them, MTO_GUI.m serves as the main file of the GUI, while MTO_CMD.m provides functionality for executing experiments via the command line.
II-C Code Pattern
II-C1 Algorithm
TABLE III: Properties and methods of algorithm base class.
Property or method Description
FE Number of fitness function evaluations
Gen Number of evolutionary generations
Best Best individual found for single-objective optimization
Result Result data contains Obj, CV, and Dec
getParameter() Get customized parameters for algorithm object
setParameter() Set customized parameters for algorithm object
notTerminated() Determine whether to terminate and update result data
Evaluation() Fitness function evaluation
run() Executing algorithm
1    label={list:algo},
2    caption={Algorithm implementation example for multi-population EA.},
3    captionpos=b,
4    float=htbp
5]
6classdef Algo_Example1_MP < Algorithm
7% <Multi-task> <Single-objective> <None/Constrained>
8
9properties % set parameter
10  para_example = 0.5;
11end
12
13methods
14function run(Algo, Prob)
15% initialize multiple populations for all tasks
16pops = Initialization(Algo, Prob, Individual);
17
18while Algo.notTerminated(Prob, pops) % main loop
19  for t = 1:Prob.T % for each task
20    % generate new offspring
21    off(index1) = Generation(pops{t});
22    % knowledge transfer
23    off(index2) = KnowledgeTransfer(pops)
24    % evaluate the fitness of offspring
25    off = Algo.Evaluation(off, Prob, t);
26    % environmental selection
27    pops{t} = Selection(pops{t}, off);
28  end
29end
30end
1    label={list:algo2},
2    caption={Algorithm implementation example for multifactorial EA.},
3    captionpos=b,
4    float=htbp
5]
6classdef Algo_Example2_MF < Algorithm
7% <Multi-task> <Multi-objective> <None/Constrained>
8
9properties % set parameter
10  para_example = 0.2;
11end
12
13methods
14function run(Algo, Prob)
15% initialize populations for multifactorial evolution
16population=Initialization_MF(Algo, Prob, Individual);
17
18while Algo.notTerminated(Prob, pops) % main loop
19  % generate new offspring with random mating
20  off = Generation(population);
21  % evaluate based on offspring skill factor
22  for t = 1:Prob.T
23    idx = [off.MFFactor] == t;
24    off(idx) = Algo.Evaluation(off(idx), Prob, t);
25  end
26  % environmental selection
27  population = Selection(population, off);
28end
29end
All algorithms within MToP inherit from the algorithm base class Algorithm.m. The properties and methods of this base class are detailed in Table III. This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself. The MTEAs implemented in MToP are outlined in Table IV. Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems. Specific algorithms belonging to this category are listed in Table V.
Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively. The algorithm labels provided in the second line serve as identifiers for classification within the GUI. The run() function within each class orchestrates algorithm execution, with Algo representing the object itself and Prob denoting the problem object to be solved. Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP.
In Listing LABEL:list:algo, the variable pops is of cell data type, housing multiple populations as popspopulation1, population2, population3…. Each individual within pops is an Individual object, containing properties such as objective value (Obj), constraint violation (CV), and decision variables (Dec). Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the population variable, directly containing individuals as population(individual1:factor1, individual2:factor1, individual3:factor2…), along with an additional property (MFFactor).
Subsequently, the primary loop commences with the invocation of the notTerminated() function, a component of the algorithm base class. This function oversees data updates and generation counting within the loop. During the main loop, distinct operations are carried out for offspring generation Generation(), offspring evaluation Evaluation(), and environmental selection Selection().
In the context of knowledge transfer, the multi-population algorithm requires the implementation of the KnowledgeTransfer() function to acquire knowledge from other tasks. Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function. The Generation() function operates on the decision variables (Dec) of offspring individuals and is tailored to specific algorithms. Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object (Prob) for offspring fitness evaluation. This entails calculating the objective value (Obj) and constraint violation (CV) for all individuals. Following each loop iteration, the environmental selection function Selection() is invoked to update the new population. While MToP offers universal Selection() functions, specific algorithms also have the option to reimplement this function. Subsequently, the code progresses to the next loop and invokes notTerminated() to document changes. The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.
TABLE IV: Multitask evolutionary algorithms included in MToP (current version).
Algorithm Year Publication Objective Task Framework Constraint Description
MFEA [19] 2016 TEVC Single Multi Multifactorial Yes Multifactorial evolution-based GA for multitask optimization
LDA-MFEA [46] 2017 CEC Single Multi Multifactorial - Linearized domain adaptation in MFEA
MFDE&MFPSO [47] 2017 CEC Single Multi Multifactorial Yes Multifactorial evolution-based DE and PSO
G-MFEA [48] 2019 TEVC Single Multi Multifactorial - Generalized MFEA with decision variable translation and shuffling
MFEA-GHS [49] 2019 ESWA Single Multi Multifactorial Yes Genetic Transform and hyper-rectangle search in MFEA
MFEA-DV [50] 2019 CEC Single Multi Multifactorial Yes Enhanced MFEA with cross-task search direction
MFEA-II [20] 2020 TEVC Single Multi Multifactorial Yes MFEA with online transfer parameter estimation
SREMTO [51] 2020 TEVC Single Multi Multifactorial Yes Self-regulated multitask framework for varying relatedness among tasks
TLTLA [52] 2020 FNS Single Multi Multifactorial Yes MFEA with two-level inter- and intra-task transfer learning
MFEA-AKT [21] 2021 TCYB Single Multi Multifactorial - MFEA with adaptive knowledge transfer via multiple crossover operator
ASCMFDE [53] 2021 TEVC Single Multi Multifactorial Yes Intertask transfer in aligned subspace for MFDE
AT-MFEA [26] 2022 TCYB Single Multi Multifactorial Yes Affine transformation-enhanced domain adaptation for MFEA
MFEA-DGD [54] 2022 TCYB Single Multi Multifactorial - MFEA based on diffusion gradient descent
IMEA [55] 2018 GECCO Single Multi Multi-population Yes Multi-population framework for multitask optimization through island model
EMEA [24] 2019 TCYB Single Multi Multi-population Yes Explicit knowledge transfer for multitask optimization via autoencoding
MFMP [22] 2020 INS Single Multi Multi-population Yes Multi-population-based adaptive DE for multitask optimization
DEORA [33] 2022 TEVC Single Multi Multi-population - Adaptive task selection for competitive multitask optimization
MKTDE [56] 2022 TEVC Single Multi Multi-population Yes Meta-knowledge transfer-based multitask DE
MTES [57] 2022 TCYB Single Multi Multi-population Yes Multitask OpenAI-ES via gradient-free evolution multitasking
BLKT-DE [58] 2023 TCYB Single Multi Multi-population Yes Multitask DE with block-level knowledge transfer
MTSRA [44] 2023 ESWA Single Multi Multi-population - Improved adaptive DE with competitive task selection
MTEA-AD [23] 2022 TEVC Single Multi/Many Multi-population Yes MTEA with adaptive knowledge transfer via anomaly detection
MTES-KG [59] 2023 TEVC Single Multi/Many Multi-population Yes Multitask ES with knowledge-guided external sampling
BoKT [30] 2022 TEVC Single Many Multi-population - Bi-objective knowledge transfer framework for many-task optimization
TRADE [60] 2023 TCYB Single Many Multi-population Yes Transferable adaptive parameter DE for many-task optimization
MTEA-SaO [61] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive solver multitask framework with implicit knowledge transfer
KR-MTEA [62] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive multi-task EA based on knowledge reuse
SBO [41] 2019 AAAI Single/Multi Many Multi-population Yes Symbiosis in biocoenosis framework for many-task optimization
MaTDE [40] 2020 TETCI Single/Multi Many Multi-population Yes Many-task DE with adaptive archive-based knowledge transfer
EMaTO-MKT [32] 2022 TEVC Single/Multi Many Multi-population - Multi-source knowledge transfer via local distribution estimation
MO-MFEA [42] 2017 TCYB Multi Multi Multifactorial Yes Multiobjective MFEA with non-dominated sort and crowding distance
MO-MFEA-II [63] 2021 TCYB Multi Multi Multifactorial Yes Cognizant Multitasking for parameter estimation in MO-MFEA
EMT-PD [64] 2021 TSMC-S Multi Multi Multifactorial Yes Two-stage adaptive knowledge transfer based on population distribution
EMT-GS [29] 2023 TEVC Multi Multi Multifactorial Yes Generative adversarial networks for knowledge transfer
MM-DE [65] 2018 CEC Multi Multi Multi-population Yes Fast memetic DE for multiobjective multitask optimization
AMT-NSGA-II [66] 2019 TCYB Multi Multi Multi-population Yes Curbing negative influences online for seamless transfer
EMT-ET [34] 2021 TCYB Multi Multi Multi-population Yes Effective knowledge transfer approach via non-dominated sort
MOMFEA-SADE [67] 2022 TCYB Multi Multi Multi-population Yes Subspace alignment and adaptive Differential Evolution
MTEA-D-DN [68] 2023 TEVC Multi Multi Multi-population - Neighborhood as a bridge for decomposition-based knowledge transfer
MTDE-MKTA [69] 2024 TEVC Multi Multi Multi-population Yes Multitask DE with multiple knowledge types and transfer adaptation
TABLE V: Single-task evolutionary algorithms for multitask optimization included in MToP (current version).
Algorithms Year Publication Objective Category Constraint Description
GA - - Single Genetic Algorithm Yes Genetic algorithm with SBX crossover and polynomial mutation
DE - - Single Differential Evolution Yes Differential evolution algorithm with DE/rand/1/bin operator
jDE [70] 2006 TEVC Single Differential Evolution Yes Self-adapting control parameters in DE
JADE [71] 2009 TEVC Single Differential Evolution Yes Adaptive DE with fast convergence performance
rank-DE [72] 2013 TCYB Single Differential Evolution Yes Ranking-based mutation operators for DE
SHADE [73] 2013 CEC Single Differential Evolution Yes Success-history based parameter adaptation for DE
jSO [74] 2017 CEC Single Differential Evolution Yes Champion of CEC 2017 numerical optimization competition via improved DE
LSHADE44 [75] 2017 CEC Single Differential Evolution Yes L-SHADE with competing strategies applied to constrained optimization
L-SHADE [76] 2014 CEC Single Differential Evolution Yes Linear population size reduction for SHADE
CAL-SHADE [77] 2017 CEC Single Differential Evolution Yes Adaptive constraint handling technique for L-SHADE
NL-SHADE-RSP [78] 2021 CEC Single Differential Evolution - Champion of CEC 2021 numerical optimization competition via improved DE
PSO [79] 1995 ICNN Single Swarm Intelligence Yes Particle swarm optimization with global and particle best update
CSO [80] 2015 TCYB Single Swarm Intelligence Yes Competitive swarm optimizer for large-scale optimization
MPA [81] 2020 ESWA Single Swarm Intelligence Yes Marine predators algorithm for numerical optimization
EO [82] 2020 KBS Single Swarm Intelligence Yes Equilibrium optimizer for numerical optimization
AO [83] 2021 CIE Single Swarm Intelligence Yes Aquila optimizer for numerical optimization
CMA-ES [84] 2001 ECJ Single Evolution Strategy Yes Evolution strategy with derandomized covariance matrix adaptation
IPOP-CMA-ES [85] 2005 CEC Single Evolution Strategy Yes Restart CMA-ES with increasing population size
sep-CMA-ES [86] 2008 PPSN Single Evolution Strategy Yes Separable CMA-ES for large-scale optimization
xNES-as [87] 2012 GECCO Single Evolution Strategy Yes Natural ES with adaptation sampling
R1-NES [88] 2013 GECCO Single Evolution Strategy Yes A Linear Time natural ES for non-separable functions
xNES&SNES [89] 2014 JMLR Single Evolution Strategy Yes ES with adaptive natural gradients to update distribution
OpenAI-ES [90] 2017 ArXiv Single Evolution Strategy Yes Parallelized ES with standard normal distribution gradients
DES [91] 2020 TEVC Single Evolution Strategy Yes Matrix-free covariance matrix adaptation ES
MTV-DE [92] 2007 EngOpt Single Constraint Handling Yes Multiple trial vectors in DE for handling constraints
ECHT [93] 2010 TEVC Single Constraint Handling Yes Ensemble of constraint handling techniques
FROFI [94] 2016 TCYB Single Constraint Handling Yes Incorporating objective function information into the feasibility rule
C2oDE [95] 2019 TSMC-S Single Constraint Handling Yes Composite DE for constrained optimization
CORCO [1] 2020 TEVC Single Constraint Handling Yes Utilizing the correlation between constraints and objective function
DeCODE [96] 2021 TSMC-S Single Constraint Handling Yes Decomposition-based multiobjective approach for constrained optimization
VMCH [97] 2022 TEVC Single Constraint Handling Yes Voting-mechanism-based ensemble of constraint handling techniques
CEDE-DRL [98] 2023 SWEC Single Constraint Handling Yes Deep reinforcement learning assisted co-evolutionary DE
CCEF-ECHT [99] 2024 TSMC-S Single Constraint Handling Yes Competitive and cooperative ensemble of constraint handling techniques
SPEA2 [100] 2001 EUROGEN Multi Non-dominated Sort Yes Improving the strength Pareto approach for multiobjective optimization
NSGA-II [101] 2002 TEVC Multi Non-dominated Sort Yes Multiobjective GA with non-dominated sort and crowding distance
NSGA-III [4] 2014 TEVC Multi Non-dominated Sort Yes Reference-point-based nondominated sorting for multiobjective optimization
MO-CMA-ES [102] 2007 ECJ Multi Evolution Strategy - Covariance matrix adaptation for multiobjective optimization
MOEA/D [3] 2007 TEVC Multi Decomposition - Decomposition-based multiobjective evolutionary algorithm
MOEA/D-DE [103] 2009 TEVC Multi Decomposition - MOEA/D for complex Pareto sets in multiobjective optimization
CCMO [2] 2021 TEVC Multi Constraint Handling Yes Coevolutionary framework for constrained multiobjective optimization
LMOCSO [104] 2020 TCYB Multi Swarm Intelligence Yes Efficient large-scale multiobjective optimization based on CSO
CMOCSO [105] 2022 TEVC Multi Swarm Intelligence Yes Competitive and cooperative swarm optimizer for multiobjective optimization
II-C2 Problem
TABLE VI: Properties and methods of problem base class.
Property or method Description
T Number of optimization tasks
N Population size for each task
M Number of objectives for all tasks
D Number of decision variable dimensions for all tasks
Fnc Fitness function for all tasks
Lb Lower bound of decision variables for all tasks
Ub Upper bound of decision variables for all tasks
maxFE Maximum fitness function evaluations
getRunParameter() Get public parameters for problem object
getParameter() Get customized parameters for problem object
setParameter() Set customized parameters for problem object
setTasks() Set optimization tasks
getOptimum() Get optimum solutions for multi-objective optimization
1    label={list:prob},
2    caption={Problem implementation example.},
3    captionpos=b,
4    float=htbp
5]
6classdef Prob_Example < Problem
7% <Multi-task> <Multi-objective> <None>
8
9methods
10function Prob = Prob_Example(varargin)
11  Prob = Prob@Problem(varargin);
12  % set default maximum function evaluations
13  Prob.maxFE = 1000 * 100;
14end
15function setTasks(Prob)
16  Prob.T = 2;
17  % task 1
18  Prob.D(1) = 10; % variable dimensions
19  Prob.M(1) = 2; % objective number
20  Prob.Fnc{1} = @func1 % fitness function
21  Prob.Lb{1} = zeros(1, 10); % lower bound
22  Prob.Ub{1} = ones(1, 10); % upper bound
23  % task 2
24  Prob.D(2) = 20; % variable dimensions
25  Prob.M(2) = 3; % objective dimensions
26  Prob.Fnc{2} = @func2 % objective function
27  Prob.Lb{2} = [0,-ones(1, 19)]; % lower bound
28  Prob.Ub{2} = [1,ones(1, 19)]; % upper bound
29end
30function optimum = getOptimum(Prob)
31  % return optimum points for each task
32  optimum{1} = optimum_matrix1;
33  optimum{2} = optimum_matrix2;
34end
35end
All problems within MToP inherit from the problem base class Problem.m. The properties and methods of Problem.m are detailed in Table VI. As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach [19]. For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$, respectively, and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is linearly mapped as follows:
$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
The dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
A simple problem implementation example is illustrated in Listing LABEL:list:prob. The maximum number of function evaluations, denoted as maxFE, for the problem can be specified within the class constructor. The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task. Within this function, the number of tasks T is first defined, followed by a detailed setup for each task. Subsequently, parameters such as decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub are set individually for each task. For problems involving multiple tasks, M and D are represented as vector type, while Fnc, Lb, and Ub are represented as cell type. Notably, the function Fnc for each task takes the decision variable Dec as input and returns objective values Obj along with the constraint violation value CV. Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP. The existing problems available in MToP are enumerated in Table VII. In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
TABLE VII: Multitask benchmark and real-world optimization problems included in MToP (current version).
Problem Case Task Dimension Objective Special Description
CEC17-MTSO [106] 9 2 50 Single - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MaTSO [107] 6 2-50 50 Single - Single-objective many-task optimization problems
WCCI20-MTSO [108] 10 2 50 Single - Complex single-objective multitask optimization problems
WCCI20-MaTSO [108] 10 2-50 50 Single - Complex single-objective many-task optimization problems
CMT [45] 9 2 Any Single Constrained Large/partial/no feasible regions interaction and different inter-task similarity
C2TOP&C4TOP [33] 51 2/4 50 Single Competitive Different optimal objective values among competitive tasks
C-CPLX [44] 20 2 50 Single Competitive Complex competitive multitask optimization problems
CEC17-MTMO [109] 9 2 50 Multi - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MTMO [107] 10 2 50 Multi - Complex multi-objective multitask optimization problems
CEC19-MaTMO [107] 10 2-50 50 Multi - Multi-objective many-task optimization problems
WCCI20-MaTMO [108] 10 2-50 50 Multi - Complex multi-objective many-task optimization problems
CEC21-MTMO [110] 10 2 50 Multi - More complex multi-objective multi optimization problems
PKACP [30] 1 Any Any Single - Planar kinematic arm control problem as many-task optimization problem
SCP [33] 1 Any 60-120 Single Competitive Sensor coverage problem as competitive multitask optimization problem
PEPVM [61] 1 3 7 Single - Parameter extraction of photovoltaic models as multitask optimization problem
OPF [59] 5 2 24/33 Single Constrained Optimal power flow as constrained multitask optimization problem
SOPM [69] 2 3 30 Multi Constrained Synchronous optimal pulse-width modulation as constrained multiobjective multitask problem
MO-OPF [69] 4 2 24/33 Multi Constrained Multiobjective optimal power flow as constrained multiobjective multitask problem
MO-SCP [61] 1 Any 60-120 Multi Competitive Multiobjective sensor coverage problem as competitive multiobjective multitask problem
II-C3 Performance Metric
1    label={list:metric},
2    caption={Metric implementation example.},
3    captionpos=b,
4    float=htbp
5]
6function result = Metric_Example(MTOData)
7% <Metric> <Multi-objective>
8
9result.Metric = ’Min’;
10
11% Data for shown in the GUI table
12result.RowName = {MTOData.Problems.Name};
13result.ColumnName = {MTOData.Algorithms.Name};
14result.TableData = CalculateTableData(MTOData);
15
16% Data for shown in the GUI convergence plot
17result.ConvergeData = CalculateConvergeData(MTOData);
18
19% Data for shown in the GUI Pareto plot
20result.ParetoData = CalculateParetoData(MTOData);
21end
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is MTOData, which is generated during experimental execution. The function returns result, comprising properties such as Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData. The Metric property can take values of either Min’ or Max’, indicating whether a smaller or larger metric value is preferable. RowName, ColumnName, and TableData are utilized to present metric results in the GUI table. On the other hand, ConvergeData and ParetoData are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table VIII.
TABLE VIII: Metrics included in MToP (current version).
Metric Objective Task Description
Obj Single All Objective value for each task
Obj (MTS) Single Multi Multitask score of Obj for all tasks
Obj (AV) Single Multi Average Obj for all tasks
Obj (CMT) Single Multi Competitive multitask Obj for all tasks
Obj (NBR) Single Multi Number of best Obj result for all tasks
CV Single All Constraint violation for each task
FR Single All Feasible rate for each task
HV Multi All Hypervolume for each task
HV (MTS) Multi Multi Multitask score of HV for all tasks
HV (CMT) Multi Multi Competitive multitask HV for all tasks
IGD Multi All Inverted generational distance for each task
IGD+ Multi All Improved IGD plus for each task
IGD+ (MTS) Multi Multi Multitask score of IGD+ for all tasks
IGD+ (CMT) Multi Multi Competitive multitask IGD+ for all tasks
Run Time All All Algorithm running time for all tasks
III Performing Experiments with MToP
(a) Test module
(b) Experiment module
Figure 7: Graphical user interface of MToP.
To launch MToP, start by running the script file mto.m located in the root directory. This action will initialize the MToP GUI interface, which is illustrated in Fig. 7.
III-A Testing Algorithms and Problems
As depicted in Fig. 7 (a), the test module interface is structured into left, center, and right sections. The left section facilitates the selection and configuration of algorithms and problems. Upon choosing the desired labels such as Task, Objective, and Special from the drop-down buttons, the options available for Algorithm and Problem are updated accordingly. Selecting specific algorithms and problems reveals their parameters in the edit box, which can be edited by clicking on them. Once parameters are adjusted, the new values are applied to the algorithm and problem objects through the setParameter() method. Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size N, maximum function evaluations maxFE, and other specific parameters.
In the center section of the MToP GUI, users can execute the test experiment and visualize the results graphically. Upon completion of the algorithm execution on the chosen problem, the metric results of the problem are displayed in the right section of the test module.
III-B Executing Comparative Experiments
The GUI of the experiment module is displayed in Fig. 7 (b), featuring a three-column layout. The left and middle columns comprise the experimental setup, while the right column displays the results.
In the first column, users can configure the basic experiment settings. The No. of Runs textbox specifies the number of independent repeat runs for each algorithm on each problem. The No. of Results textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations. The Parallel button toggles parallel processing, enabling multiple threads to run independent experiments simultaneously. The drop-down buttons Task, Objective, and Special filter algorithms and problems based on predefined labels, similar to the test module. Users can add selected algorithms and problems by clicking the Add button after choosing them from the respective drop-down boxes at the bottom.
In the second column, the names and parameters of the added algorithms and problems can be edited by clicking on them. Once all configurations are set, clicking the green Start button initiates the experiment execution. Users can also pause or stop the experiment using the Pause and Stop buttons if needed. Once the experiment is completed, the table’s contents become available in the results area on the right part of the experiment module. Users can read and save the original data of the experiment using the Load Data and Save Data buttons.
In the results area, the first function row contains buttons for operating on the experimental data. The first button, labeled Save Table, exports the table results to xlsx, csv, and tex files. The subsequent buttons labeled Converge control the convergence y-lim types and generate convergence plots based on the selected results from the table. The Pareto button draws the Pareto front and non-dominated solutions based on the selected results.
The edit box controls the formatting type of the table result. Following this, the drop-down box allows users to select performance metrics, and subsequent drop-down boxes determine the result display type, statistical test method, base algorithm, and result highlighting type, respectively. After selecting the statistical test method, calculated statistical results denoted by “+/-/=” followed by each metric result can also be displayed in the table. MToP offers two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests. The base algorithm drop-down button controls the base algorithm in the statistical test, while the highlight button highlights the best and worst algorithm’s metric results for each problem or task.
III-C Executing via Alternative Command Line
In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line. By inputting the algorithm list, problem list, and other parameters into the script file mto.m, experiments can be executed directly from the command line. Upon completion, the experiment data MTOData.mat is returned. Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
IV Extending MToP
Users can seamlessly integrate new algorithms, problems, and performance metrics into MToP. By adhering to the code pattern outlined in Section II-C, users can develop and implement their code files with ease.
IV-A Adding New Algorithms
To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the Algorithms/ folder. This class file should inherit from the base class Algorithm.m. Subsequently, users can implement the algorithmic framework following the provided code pattern, as demonstrated in Listing LABEL:list:algo. Utilizing functions such as Initialization(), Generation(), and Selection() provided by MToP is recommended. However, users can also choose to reimplement these functions to suit their specific requirements. Additionally, users seeking to create customizable parameters that can be modified via the MToP GUI should implement the getParameter() and setParameter() APIs within the new algorithm class file.
IV-B Adding New Problems
To integrate new problems into MToP, users should create a dedicated problem class file within the Problems/ folder. This class file should inherit from the base class Problem.m. It’s advisable to define the suggested maximum function evaluations (maxFE) within the constructor of the specific problem. Subsequently, users can refer to Listing LABEL:list:prob for guidance on configuring each optimization task within the setTasks() function. In cases where optimization tasks involve multi-objective optimization problems, users need to additionally implement the getOptimum() function to provide the optimal solution set or reference points. Lastly, for setting customizable parameters that can be modified within the GUI, users must also implement the getParameter() and setParameter() functions.
IV-C Adding New Performance Metrics
To incorporate new performance metrics into MToP, users should create a new metrics function file within the Metrics/ folder. Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object MTOData and return the calculated result object. For the metrics displayed in the GUI table, ensure that the RowName, ColumnName, and TableData fields are properly populated. Typically, the RowName for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task. The ColumnName typically represents the algorithm name. Implement the calculation function for the specific performance metric in the TableData field. For plotting metric convergence, ensure the ConvergeData field is populated. In the case of multi-objective optimization, include non-dominated solutions in the ParetoData field.
V Conclusion and Future Work
This paper introduces MToP, an open-source MATLAB platform designed for EMT. MToP features a user-friendly GUI, a rich collection of algorithms and problems, and convenient code structures. The current version of MToP contains over 40 MTEAs, more than 40 single-task EAs capable of handling MTO problems, over 150 benchmark MTO problems, and several real-world applications of EMT. This paper provides detailed descriptions of MToP’s functional modules, file structure, and code patterns, as well as guidelines for conducting experiments using MToP.
While MToP has undergone careful reimplementation, modification, and testing of included algorithms and problems, it may still contain some reproducing errors and bugs. Continuous efforts are underway to identify and rectify such issues, with updates regularly posted on GitHub. The platform encourages feedback and contributions from followers and researchers, with many codes already received and integrated from the community.
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of multitask combinatorial optimization problems [5, 6] and more real-world applications [8, 111] in the future.
Acknowledgement
All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles. The code provided by the original authors undergoes modifications based on the source code during its integration into MToP.
The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO [36], for which the authors are also thankful.