### v3

# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

Abstract
Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past years. It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques. Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 20 performance metrics. Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 40 popular single-task evolutionary algorithms to address MTO problems. MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization. More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
Index Terms:
Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm.
I Introduction
Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency. Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and user-friendly nature. Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6]. In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking [7]. A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ optimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
$\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,F_{K}(\bm{x}_{K})\Big{]},\\ \textit{s.t. }\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$ (1)
where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective optimization task, the $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task. For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space [8] as follows:
$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. EMT has found successful applications in various domains, including engineering scheduling [5, 9], nonlinear equation systems [10], feature selection in machine learning [11], anomaly detection [12], point cloud registration [13, 14], and reinforcement learning [15].
To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques [16, 17]. Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources [18, 19]. The first attempt of EMT can be traced back to the multifactorial EA [8], which introduced an implicit knowledge representation via random mating among optimization tasks. Subsequently, numerous EAs tailored for MTO have emerged. These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks [20, 21], or a multi-population framework allocating separate populations for each task [22, 23]. Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space [8, 21], autoencoding [24, 25], affine transformation [26, 27, 28], and adversarial generative models [29] have been proposed in MTEAs. Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment [20, 30, 31], knowledge selection [23, 32], and historical feedback [33, 34, 35], have also been investigated. Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes. This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs. Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field. Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field. In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization. For instance, PlatEMO [36] is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems. EvoX [37] is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training. Evo-Panel [38] created a comprehensive dynamic visualization tool to illustrate the detailed procedures of different optimization algorithms in solving numerical benchmark functions. EDOLAB [39] is an optimization tool specifically for dynamic optimization.
However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks. Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics. Specifically, single-task metrics evaluate the performance of each task in the MTO problem, while multitask metrics assess the overall performance of the MTO problem. These bring uncertainty to the pattern of results display and analysis. These distinctive requirements present significant challenges for the implementation of MTEAs in existing tools and platforms. Moreover, EMT has been extended to the subfields of many-task optimization [40, 41], multi-objective multitask optimization [42, 43], competitive multitask optimization [33, 44], and constrained multi-task optimization [45]. Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. MToP is designed to provide a comprehensive software platform for researchers to evaluate MTEAs on benchmark MTO problems and explore real-world applications. The reason for using MATLAB is that it is a widely used programming language in the field of EC. Moreover, MATLAB provides a user-friendly graphical interface and a powerful numerical computing environment, making it an ideal choice for developing MToP. More importantly, the existing MTEAs and MTO problems are mainly implemented in MATLAB, which can be directly integrated into MToP. The main contributions of MToP are summarized as follows:
  1. 1.
    MToP features a user-friendly graphical user interface (GUI) comprising test, experiment, and data-process modules. These modules facilitate researchers in understanding problem characteristics, conducting comparative experiments, solving problems in parallel, statistically analyzing results, plotting result figures, and managing experimental data. Moreover, MToP offers a modular implementation of algorithms, problems and performance metrics. On top of these, it provides an extensive public application programming interface (API) with template functions for population initialization, function evaluation, evolutionary operators, and environmental selection.
  2. 2.
    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API. Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types. Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types. In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications. Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities.
  3. 3.
    MToP is designed to be easily extended, allowing for the seamless addition of new algorithms, problems, and metrics. By adhering to established coding patterns and implementing functionality based on the public API, new code can be seamlessly integrated and utilized within MToP. Given its status as a completely open-source project, researchers have the opportunity to leverage existing algorithms, problems, and metrics as the foundation for novel ideas. Through the collaborative platform GitHub, MToP undergoes continuous updates and enhancements, ensuring that it remains at the forefront of EMT research [^1] MToP has been continuously updated since Sep. 21, 2021, when it was uploaded to GitHub at https://github.com/intLyc/MTO-Platform..
The rest of this paper is organized as follows. Section II elaborates on the architecture of MToP. Section III describes approaches for performing experiments with MToP. Section IV introduces how to extend the contents of MToP. Finally, the conclusion and future work are given in Section V.
II Architecture of MToP
MToP consists of key modules including the test, experiment, and data processing modules, which facilitate algorithm execution, visualization, and data analysis. These components allow for flexible experimentation through both GUI and command line interfaces. Following standardized code patterns, MToP integrates various algorithms and performance metrics.
II-A Functional Modules
II-A1 Test Module
(a) Function landscapes (1D)
(b) Function landscapes (2D)
(c) Feasible regions
(d) Pareto front
(e) Metric convergence
Figure 1: Examples of graphical display in the test module of MToP. (a) and (b) illustrate the landscapes of single-objective problems with different tasks in the one- and two-dimensional unified search space, respectively. (c) shows the feasible regions of a single-objective problem with different tasks in the two-dimensional unified search space. (d) depicts the Pareto front of a multi-objective problem with multiple tasks. (e) displays the convergence behavior of metrics after executing algorithms on problems.
The test module is designed to assist researchers in analyzing the characteristics of MTO problems and algorithms. It facilitates running the selected algorithm on the chosen problem within MToP. The module includes visualization functions for both single-objective and multi-objective MTO problems. Users can select different options to display depending on the problem type, resulting in different visual outputs, as demonstrated below. In the case of single-objective MTO problems, the landscapes of the objective function with different tasks in the one- and two-dimensional unified search space can be depicted, as shown in Fig. 1 (a)-(b). For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be visualized, as illustrated in Fig. 1 (c). Regarding multi-objective MTO problems, the Pareto front figures can be generated for multiple tasks, as depicted in Fig. 1 (d). Furthermore, the convergence behavior of metrics can be plotted after executing algorithms on problems, as exemplified in Fig. 1 (e).
In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior. Figure 2 illustrates an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space. CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives. The optimal Pareto set of the first task contains diversity dimension $1$ with optimal value range $[0,1]$ and convergence dimension $[2,50]$ with optimal value $0.5$. The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.4$. Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
(a) Gen=1
(b) Gen=100
(c) Gen=200
(d) Gen=500
(e) Gen=700
(f) Gen=1000
(g) Gen=1
(h) Gen=100
(i) Gen=200
(j) Gen=500
(k) Gen=700
(l) Gen=1000
Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space and decision space with the number of generations. The top row (a)-(f) shows the population in the objective space, while the bottom row (g)-(l) displays the population in the decision space.
In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space. As evolution progresses through the first and middle phases (Gen=100 to 200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity. Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity. By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task. The schematics indicate that the population for the second task occupies a similar position in the decision space as the first task. However, these two tasks do not precisely overlap in the decision space. This observation suggests the presence of negative knowledge transfer, which could be contributing to this discrepancy. In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
Figure 3: The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
The workflow of API calls for executing an algorithm on a problem is depicted in Fig. 3. The process initiates with the <Start_Button> event in <MTO_GUI.m>, triggering the parameter setting API SetParameter() within the selected algorithm <MFEA.m>. Subsequently, it proceeds to the Run() function of the algorithm, which initially invokes the Initialization() API of the algorithm base class <Algorithm.m>, and then iterates through the function evaluation API Evaluation(). During the initialization phase, the parameter settings of the problem <CEC17_MTP1.m> to be solved are obtained via the problem base class <Problem.m>. Similarly, the Evaluation() function retrieves the specific optimization task object through the problem base class. Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function <Obj.m> via the CallMetric() API.
All functions within MToP that execute algorithms, including subsequent experiment modules, adhere to this procedure. Additionally, both algorithm and problem parameters can be directly modified through the GUI of MToP.
II-A2 Experiment Module
TABLE I: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
GA MFEA MTEA-AD
CI-HS-T1 9.0426e-01 (6.16e-02) 8.4445e-01 (5.51e-02) + 9.3438e-01 (5.03e-02) =
CI-HS-T2 4.2400e+02 (4.90e+01) 2.8119e+02 (5.09e+01) + 3.1099e+02 (4.11e+01) +
CI-MS-T1 5.4562e+00 (1.52e+00) 5.8307e+00 (8.23e-01) - 4.5085e+00 (2.69e-01) +
CI-MS-T2 4.2474e+02 (6.45e+01) 3.1448e+02 (4.41e+01) + 3.1660e+02 (4.39e+01) +
CI-LS-T1 2.1167e+01 (1.82e-01) 2.0234e+01 (8.35e-02) + 2.1086e+01 (2.82e-01) =
CI-LS-T2 4.1476e+03 (6.34e+02) 4.3638e+03 (4.88e+02) = 4.5958e+03 (6.64e+02) -
PI-HS-T1 4.2186e+02 (5.27e+01) 6.6221e+02 (1.11e+02) - 4.4662e+02 (4.13e+01) -
PI-HS-T2 8.6994e+01 (1.93e+01) 8.9970e+01 (1.67e+01) = 1.2177e+02 (2.88e+01) -
PI-MS-T1 5.4736e+00 (1.42e+00) 4.9276e+00 (4.76e-01) = 4.7264e+00 (3.50e-01) +
PI-MS-T2 2.4152e+04 (1.00e+04) 1.0488e+04 (4.77e+03) + 9.1688e+03 (4.69e+03) +
PI-LS-T1 5.6540e+00 (1.88e+00) 1.9583e+01 (2.71e+00) - 5.3895e+00 (9.11e-01) =
PI-LS-T2 1.4062e+01 (2.48e+00) 1.9371e+01 (3.26e+00) - 5.1076e+00 (1.44e+00) +
NI-HS-T1 2.9452e+04 (1.49e+04) 9.7165e+03 (4.93e+03) + 1.3459e+04 (4.77e+03) +
NI-HS-T2 4.5183e+02 (6.26e+01) 3.7757e+02 (7.84e+01) + 3.8559e+02 (6.26e+01) +
NI-MS-T1 9.0462e-01 (6.19e-02) 9.3064e-01 (5.16e-02) = 9.8876e-01 (4.48e-02) -
NI-MS-T2 3.8662e+01 (3.97e+00) 2.7501e+01 (3.26e+00) + 2.4831e+01 (3.76e+00) +
NI-LS-T1 4.2895e+02 (5.41e+01) 7.4411e+02 (1.30e+02) - 4.6340e+02 (4.97e+01) -
NI-LS-T2 4.3308e+03 (5.19e+02) 4.5452e+03 (4.90e+02) = 4.7449e+03 (6.45e+02) -
+ / - / = Base 8 / 5 / 5 9 / 6 / 3
(a) Convergence
(b) Pareto front
Figure 4: schematic plotting by the experiment module of MToP. (a) Convergence plot of algorithms on problems. (b) Pareto front plot of multi-objective problems.
The experiment module offers functionalities for conducting comprehensive experiments and analyzing results. Users can execute multiple independent runs of various algorithms to tackle multiple MTO problems. Upon completion of the runs, data including experimental settings, objective values, and decision variables are recorded and stored. These data serve as the basis for generating metric results, conducting significance tests, and creating schematic plots.
Metrics functions operating on the data are categorized into single-task and multitask metrics. The selection of single-task metrics presents the results of each task of the MTO problem in tabular format within MToP while choosing to multitask metrics displays the results of each MTO problem. Significance test results are also included in the table alongside the metric results. All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table I.
Convergence plots are generated based on the selected table data to facilitate the comparison of convergence behaviors among different algorithms, as depicted in Fig. 4 (a). The convergence plot function can be applied to both single- and multi-objective problems. Furthermore, for multi-objective MTO algorithms and problems, the Pareto front can be plotted, as illustrated in Fig. 4 (b).
II-A3 Data Process Module
TABLE II: Properties of experimental data.
Property Description
Reps Repetitions number of independent runs
Algorithms Algorithms data contains names and parameter settings
Problems Problems data contains names and parameter settings
Results Total results data contains Obj, CV, and Dec
RunTimes Running time data of algorithms on problems
Metrics Calculated metric results data
Each unique execution within the experiment module is considered a data object and is saved to a data file named <MTOData.mat>. The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table II. This saved data can subsequently be reloaded by the experiment module of MToP.
With the data process module, the data can be merged or split, enabling data reuse and customization. The processing of data is categorized into three types: independent runs, algorithms, and problems. If several sets of data with fewer independent runs are available, the data can be merged to create new data with a larger number of independent runs. Similarly, in cases where existing data needs augmentation with new algorithms for comparison, it is unnecessary to re-execute all comparative experiments. Instead, only the new algorithms need to be executed, and the resulting data can be merged with the original data. Likewise, if not all algorithms within the data are required, the data can be split and re-merged accordingly. The data processing for problems follows a similar approach, wherein new problems can be added or deleted as needed.
In practical use, the specific operations for splitting and merging data are as follows:
  • Data Splitting: If an experiment involves running both the MFEA and EMEA algorithms on same problems, the data can be split by selecting it in the <Data Tree> and clicking the <Algorithms Split> button to generate separate datasets for each algorithm. Similarly, data can also be split by independent runs or problems using the respective options.
  • Data Merging: If the MFEA and EMEA algorithms were run on same problems in separate experiments, the data can be merged by selecting both datasets and clicking the <Algorithms Merge> button, provided the settings are identical except for the algorithm. This process can also be used to merge datasets by independent runs or problems.
II-B File Structure
Figure 5: File structure of MToP. The root directory contains the main script file <mto.m> and four subfolders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
In the root directory <MTO/> of MToP, you’ll find a script file named <mto.m> and four folders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>. The structure of these folders is depicted in Fig. 5.
The <Algorithms/> folder contains various categories of algorithms, utility functions, and the algorithm base class <Algorithm.m>. Specific algorithm files such as <MFEA.m>, <MO_MFEA.m>, and <GA.m> are stored within their respective classification folders. The <Algorithms/Utils/> folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
The <Problems/> folder houses various categories of problems, the base function folder, and the problem base class <Problem.m>. Specific problem files like <WCCI20_MTSO1.m>, <CMT3.m>, and <CEC19_MaTMO6.m> are stored under their corresponding classification folders. Real-world application problem files are located in the <Real-World Application/> folder, while the <Problems/Base/> folder contains public base functions for problems.
The <Metrics/> folder contains all result evaluation metrics, organized into subfolders for single-objective <Metrics/Single-objective/> and multi-objective <Metrics/Multi-objective> optimization. Specific metric files include <Obj.m>, <Obj_MTS.m>, and <IGD.m>, among others.
Lastly, the <GUI/> folder houses all files used by the GUI of MToP. Among them, <MTO_GUI.m> serves as the main file of the GUI, while <MTO_CMD.m> provides functionality for executing experiments via the command line.
II-C Code Pattern
II-C1 Algorithm
TABLE III: Properties and methods of algorithm base class.
Property or method Description
FE Number of fitness function evaluations
Gen Number of evolutionary generations
Best Best individual found for single-objective optimization
Result Result data contains Obj, CV, and Dec
getParameter() Get customized parameters for algorithm object
setParameter() Set customized parameters for algorithm object
notTerminated() Determine whether to terminate and update result data
Evaluation() Fitness function evaluation
run() Executing algorithm
1 label={list:algo},
2 caption={Algorithm implementation example for multi-population EA.},
3 captionpos=b,
4 float=htbp
5]
6classdef Algo_Example1_MP < Algorithm
7% <Multi-task> <Single-objective> <None/Constrained>
8
9properties % set parameter
10 para_example = 0.5;
11end
12
13methods
14function run(Algo, Prob)
15% initialize multiple populations for all tasks
16pops = Initialization(Algo, Prob, Individual);
17
18while Algo.notTerminated(Prob, pops) % main loop
19 for t = 1:Prob.T % for each task
20 % generate new offspring
21 off(index1) = Generation(pops{t});
22 % knowledge transfer
23 off(index2) = KnowledgeTransfer(pops)
24 % evaluate the fitness of offspring
25 off = Algo.Evaluation(off, Prob, t);
26 % environmental selection
27 pops{t} = Selection(pops{t}, off);
28 end
29end
30end
1 label={list:algo2},
2 caption={Algorithm implementation example for multifactorial EA.},
3 captionpos=b,
4 float=htbp
5]
6classdef Algo_Example2_MF < Algorithm
7% <Multi-task> <Multi-objective> <None/Constrained>
8
9properties % set parameter
10 para_example = 0.2;
11end
12
13methods
14function run(Algo, Prob)
15% initialize populations for multifactorial evolution
16population=Initialization_MF(Algo, Prob, Individual);
17
18while Algo.notTerminated(Prob, pops) % main loop
19 % generate new offspring with random mating
20 off = Generation(population);
21 % evaluate based on offspring skill factor
22 for t = 1:Prob.T
23 idx = [off.MFFactor] == t;
24 off(idx) = Algo.Evaluation(off(idx), Prob, t);
25 end
26 % environmental selection
27 population = Selection(population, off);
28end
29end
All algorithms within MToP inherit from the algorithm base class <Algorithm.m>. The properties and methods of this base class are detailed in Table III. This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself. The MTEAs implemented in MToP are outlined in Table LABEL:supp:tab:algo-mt of the supplementary files. Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems. Specific algorithms belonging to this category are listed in Table LABEL:supp:tab:algo-st.
Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively. The algorithm labels provided in the second line serve as identifiers for classification within the GUI. The run() function within each class orchestrates algorithm execution, with <Algo> representing the object itself and <Prob> denoting the problem object to be solved. Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP.
In Listing LABEL:list:algo, the variable <pops> is of <cell> data type, housing multiple populations as <population1, population2, population3...>. Each individual within <pops> is an <Individual> object, containing properties such as objective value <Obj>, constraint violation <CV>, and decision variables <Dec>. Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the <population> variable, directly containing individuals as <individual1:factor1, individual2:factor1, individual3:factor2...>, along with an additional property <MFFactor>.
Subsequently, the primary loop commences with the invocation of the notTerminated() function, a component of the algorithm base class. This function oversees data updates and generation counting within the loop. During the main loop, distinct operations are carried out for offspring generation Generation(), offspring evaluation Evaluation(), and environmental selection Selection().
In the context of knowledge transfer, the multi-population algorithm requires the implementation of the KnowledgeTransfer() function to acquire knowledge from other tasks. Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function. The Generation() function operates on the decision variables <Dec> of offspring individuals and is tailored to specific algorithms. Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object <Prob> for offspring fitness evaluation. This entails calculating the objective value <Obj> and constraint violation <CV> for all individuals. Following each loop iteration, the environmental selection function Selection() is invoked to update the new population. While MToP offers universal Selection() functions, specific algorithms also have the option to reimplement this function. Subsequently, the code progresses to the next loop and invokes notTerminated() to document changes. The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.
II-C2 Problem
TABLE IV: Properties and methods of problem base class.
Property or method Description
T Number of optimization tasks
N Population size for each task
M Number of objectives for all tasks
D Number of decision variable dimensions for all tasks
Fnc Fitness function for all tasks
Lb Lower bound of decision variables for all tasks
Ub Upper bound of decision variables for all tasks
maxFE Maximum fitness function evaluations
getRunParameter() Get public parameters for problem object
getParameter() Get customized parameters for problem object
setParameter() Set customized parameters for problem object
setTasks() Set optimization tasks
getOptimum() Get optimal solutions for multi-objective optimization
1 label={list:prob},
2 caption={Problem implementation example.},
3 captionpos=b,
4 float=htbp
5]
6classdef Prob_Example < Problem
7% <Multi-task> <Multi-objective> <None>
8
9methods
10function Prob = Prob_Example(varargin)
11 Prob = Prob@Problem(varargin);
12 % set default maximum function evaluations
13 Prob.maxFE = 1000 * 100;
14end
15function setTasks(Prob)
16 Prob.T = 2;
17 % task 1
18 Prob.D(1) = 10; % variable dimensions
19 Prob.M(1) = 2; % objective number
20 Prob.Fnc{1} = @func1 % fitness function
21 Prob.Lb{1} = zeros(1, 10); % lower bound
22 Prob.Ub{1} = ones(1, 10); % upper bound
23 % task 2
24 Prob.D(2) = 20; % variable dimensions
25 Prob.M(2) = 3; % objective dimensions
26 Prob.Fnc{2} = @func2 % objective function
27 Prob.Lb{2} = [0,-ones(1, 19)]; % lower bound
28 Prob.Ub{2} = [1,ones(1, 19)]; % upper bound
29end
30function optimum = getOptimum(Prob)
31 % return optimum points for each task
32 optimum{1} = optimum_matrix1;
33 optimum{2} = optimum_matrix2;
34end
35end
All problems within MToP inherit from the problem base class <Problem.m>. The properties and methods of <Problem.m> are detailed in Table IV. As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach [8]. It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
A simple problem implementation example is illustrated in Listing LABEL:list:prob. The maximum number of function evaluations, denoted as <maxFE>, for the problem can be specified within the class constructor. The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task. Within this function, the number of tasks <T> is first defined, followed by a detailed setup for each task. Subsequently, parameters such as decision variable dimensions <D>, objective dimensions <M>, fitness function <Fnc>, lower bound <Lb>, and upper bound <Ub> are set individually for each task. For problems involving multiple tasks, <M> and <D> are represented as <vector> type, while <Fnc>, <Lb>, and <Ub> are represented as <cell> type. Notably, the function <Fnc> for each task takes the decision variable <Dec> as input and returns objective values <Obj> along with the constraint violation value <CV>. Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP. The existing problems available in MToP are enumerated in Table LABEL:supp:tab:prob-mt of the supplementary files. In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
II-C3 Performance Metric
1 label={list:metric},
2 caption={Metric implementation example.},
3 captionpos=b,
4 float=htbp
5]
6function result = Metric_Example(MTOData)
7% <Metric> <Multi-objective>
8
9result.Metric = ’Min’;
10
11% Data for shown in the GUI table
12result.RowName = {MTOData.Problems.Name};
13result.ColumnName = {MTOData.Algorithms.Name};
14result.TableData = CalculateTableData(MTOData);
15
16% Data for shown in the GUI convergence plot
17result.ConvergeData = CalculateConvergeData(MTOData);
18
19% Data for shown in the GUI Pareto plot
20result.ParetoData = CalculateParetoData(MTOData);
21end
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is <MTOData>, which is generated during experimental execution. The function returns <result>, comprising properties such as <Metric>, <RowName>, <ColumnName>, <TableData>, <ConvergeData>, and <ParetoData>. The <Metric> property can take values of either ’Min’ or ’Max’, indicating whether a smaller or larger metric value is preferable. <RowName>, <ColumnName>, and <TableData> are utilized to present metric results in the GUI table. On the other hand, <ConvergeData> and <ParetoData> are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table V.
TABLE V: Metrics included in MToP. The table lists the metric name, optimization objective, task type, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. ST and MT represent single-task and multi-task, respectively
Metric Objective Task Description
Obj SO ST/MT Objective value for each task
Obj (AV) SO MT Average Obj for all tasks
Obj (UV) SO MT Unified average Obj for all tasks
Obj (MTS) SO MT Multitask score of Obj for all tasks
Obj (CMT) SO MT Competitive multitask Obj for all tasks
Obj (NBR) SO MT Number of best Obj result for all tasks
CV SO ST/MT Constraint violation for each task
FR SO ST/MT Feasible rate for each task
HV MO ST/MT Hypervolume for each task
HV (MTS) MO MT Multitask score of HV for all tasks
HV (CMT) MO MT Competitive multitask HV for all tasks
IGD MO ST/MT Inverted generational distance for each task
IGD (AV) MO MT Average IGD for all tasks
IGD+ (MTS) MO MT Multitask score of IGD for all tasks
IGD (CMT) MO MT Competitive multitask IGD for all tasks
IGD+ MO ST/MT Improved IGD plus for each task
IGD+ (MTS) MO MT Multitask score of IGD+ for all tasks
IGD+ (CMT) MO MT Competitive multitask IGD+ for all tasks
Spread MO ST/MT Spread metric for each task
Spread (CMT) MO MT Competitive multitask spread for all tasks
Run Time SO/MO ST/MT Algorithm running time for all tasks
III Performing Experiments with MToP
(a) Test Module
(b) Experiment Module
Figure 6: Graphical user interface of MToP. The test module (a) is used for testing algorithms and problems. The graphical display is shown in the center section of the test module. The experiment module (b) is used for executing comparative experiments. The results of the experiment are displayed in the right section of the experiment module.
To launch MToP, start by running the script file <mto.m> located in the root directory. This action will initialize the MToP GUI interface, which is illustrated in Fig. 6. The GUI interface of MToP requires MATLAB R2020b or later versions to run, while command-line execution can be done with any version. In addition, if parallel execution is required, MATLAB Parallel Computing Toolbox is needed.
III-A Testing Algorithms and Problems
As depicted in Fig. 6 (a), the test module interface is structured into left, center, and right sections. The left section facilitates the selection and configuration of algorithms and problems. Upon choosing the desired labels such as <Task>, <Objective>, and <Special> from the drop-down buttons, the options available for <Algorithm> and <Problem> are updated accordingly. Selecting specific algorithms and problems reveals their parameters in the edit box, which can be edited by clicking on them. Once parameters are adjusted, the new values are applied to the algorithm and problem objects through the setParameter() method. Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size <N>, maximum function evaluations <maxFE>, and other specific parameters.
In the center section of the MToP GUI, users can execute the test experiment and visualize the results graphically. Upon completion of the algorithm execution on the chosen problem, the metric results of the problem are displayed in the right section of the test module.
III-B Executing Comparative Experiments
The GUI of the experiment module is displayed in Fig. 6 (b), featuring a three-column layout. The left and middle columns comprise the experimental setup, while the right column displays the results.
In the first column, users can configure the basic experiment settings. The <No. of Runs> textbox specifies the number of independent repeat runs for each algorithm on each problem. The <No. of Results> textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations. The <Parallel> button toggles parallel processing, allowing independent experiments to run simultaneously across multiple process. The number of concurrent experiments depends on the number of processor cores. If the number of parallel experiments exceeds the available cores, MATLAB automatically handles the scheduling. For multiple algorithms and problems, each algorithm processes a problem in parallel. After one algorithm finishes, the next starts, repeating until all algorithms have processed the problem, then moving to the next problem. Hence, the number of algorithms and problems is not inherently limited. The drop-down buttons <Task>, <Objective>, and <Special> filter algorithms and problems based on predefined labels, similar to the test module. Users can add selected algorithms and problems by clicking the <Add> button after choosing them from the respective drop-down boxes at the bottom.
In the second column, the names and parameters of the added algorithms and problems can be edited by clicking on them. Once all configurations are set, clicking the green <Start> button initiates the experiment execution. Users can also pause or stop the experiment using the <Pause> and <Stop> buttons if needed. Once the experiment is completed, the table’s contents become available in the results area on the right part of the experiment module. Users can read and save the original data of the experiment using the <Load Data> and <Save Data> buttons.
In the results area, the first function row contains buttons for operating on the experimental data. The first button, labeled <Save Table>, exports the table results to xlsx, csv, and tex files. The subsequent buttons labeled <Converge> control the convergence y-lim types and generate convergence plots based on the selected results from the table. The <Pareto> button draws the Pareto front and non-dominated solutions based on the selected results.
The edit box controls the formatting type of the table result. Following this, the drop-down box allows users to select performance metrics, and subsequent drop-down boxes determine the result display type, statistical test method, base algorithm, and result highlighting type, respectively. After selecting the statistical test method, calculated statistical results denoted by “+ / - / =” followed by each metric result can also be displayed in the table. MToP offers two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests. The base algorithm drop-down button controls the base algorithm in the statistical test, while the highlight button highlights the best and worst algorithm’s metric results for each problem or task.
III-C Executing via Alternative Command Line
In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line. By inputting the algorithm list, problem list, and other parameters into the script file <mto.m>, experiments can be executed directly from the command line. For example, running the command mto({MFEA, EMEA}, {CMT1, CMT2}, 30) will execute the MFEA and EMEA algorithms on the CMT1 and CMT2 problems independently 30 times. Upon completion, the experiment data <MTOData.mat> is returned. Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
IV Extending MToP
Users can seamlessly integrate new algorithms, problems, and performance metrics into MToP. By adhering to the code pattern outlined in Section II-C, users can develop and implement their code files with ease.
IV-A Adding New Algorithms
To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the <Algorithms/> folder. This class file should inherit from the base class <Algorithm.m>. Subsequently, users can implement the algorithmic framework following the provided code pattern, as demonstrated in Listing LABEL:list:algo. Utilizing functions such as Initialization(), Generation(), and Selection() provided by MToP is recommended. However, users can also choose to reimplement these functions to suit their specific requirements. Additionally, users seeking to create customizable parameters that can be modified via the MToP GUI should implement the getParameter() and setParameter() APIs within the new algorithm class file.
IV-B Adding New Problems
To integrate new problems into MToP, users should create a dedicated problem class file within the <Problems/> folder. This class file should inherit from the base class <Problem.m>. It’s advisable to define the suggested maximum function evaluations <maxFE> within the constructor of the specific problem. Subsequently, users can refer to Listing LABEL:list:prob for guidance on configuring each optimization task within the setTasks() function. In cases where optimization tasks involve multi-objective optimization problems, users need to additionally implement the getOptimum() function to provide the optimal solution set or reference points. Lastly, for setting customizable parameters that can be modified within the GUI, users must also implement the getParameter() and setParameter() functions.
IV-C Adding New Performance Metrics
To incorporate new performance metrics into MToP, users should create a new metrics function file within the <Metrics/> folder. Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object <MTOData> and return the calculated <result> object. For the metrics displayed in the GUI table, ensure that the <RowName>, <ColumnName>, and <TableData> fields are properly populated. Typically, the <RowName> for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task. The <ColumnName> typically represents the algorithm name. Implement the calculation function for the specific performance metric in the <TableData> field. For plotting metric convergence, ensure the <ConvergeData> field is populated. In the case of multi-objective optimization, include non-dominated solutions in the <ParetoData> field.
IV-D Extending Optimization Types
There are also some areas, such as sequential transfer optimization and multiform optimization, in the evolutionary transfer optimization field similar to evolutionary multitasking [18]. Although MToP is currently focused on evolutionary multitasking, it is possible to extend the platform to include these new optimization types. Users can refer to the existing code patterns in MToP to develop new optimization types. Specifically, for the sequential transfer optimization, it is possible to set the target task as the first task and the remaining tasks as source tasks for algorithms and problems. The metric calculation can be adjusted to evaluate the performance of the target task. For multiform optimization, it can dynamically construct multiform tasks within the implemented algorithms depending on the problem to be solved.
V Conclusion and Future Work
This paper introduces MToP, an open-source MATLAB platform designed for EMT. MToP features a user-friendly GUI, a rich collection of algorithms and problems, and convenient code structures. The current version of MToP contains over 40 MTEAs, more than 40 single-task EAs capable of handling MTO problems, over 150 benchmark MTO problems, and several real-world applications of EMT. This paper provides detailed descriptions of MToP’s functional modules, file structure, and code patterns, as well as guidelines for conducting experiments using MToP.
While MToP has undergone careful reimplementation, modification, and testing of included algorithms and problems, it may still contain some reproducing errors and bugs. Continuous efforts are underway to identify and rectify such issues, with updates regularly posted on GitHub. The platform encourages feedback and contributions from followers and researchers, with many codes already received and integrated from the community.
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of many-objective optimization codes [46, 4], multitask combinatorial optimization codes [5, 6] and more real-world applications [9, 47] in the future.
Acknowledgement
All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles. The code provided by the original authors undergoes modifications based on the source code during its integration into MToP.
The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO [36], for which the authors are also thankful.